### v1

# AQORA: A Learned Adaptive Query Optimizer for Spark SQL

Abstract
Recent studies have identified two main approaches to improve query optimization: learned query optimization (LQO), which generates or selects better query plans before execution based on models trained in advance, and adaptive query processing (AQP), which adapts the query plan during execution based on statistical feedback collected at runtime. Although both approaches have shown promise, they also face critical limitations. LQO must commit to a fixed plan without access to actual cardinalities and typically rely on a single end-to-end feedback signal, making learning inefficient. On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience. In this paper, we present AQORA, an adaptive query optimizer with a reinforcement learning architecture that combines the strengths of both LQO and AQP. AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration. Experiments show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
I Introduction
Query optimization plays a central role in database systems, determining the physical plan for a given query. Typically, this process takes place prior to query execution. The query optimizer selects a fixed execution plan based on estimated data statistics, analytical cost models, and a predefined set of transformation rules. Once execution begins, the query processing engine adheres strictly to the chosen plan, without the opportunity to revise it based on actual runtime behavior.
In response to the limitations of static, pre-execution query planning, subsequent research has evolved along two primary directions. One line of work focuses on developing learning-based algorithms that assist or replace the traditional query optimizer in making key planning decisions (such as join ordering, operator selection, or cardinality estimation) prior to execution beginning[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. These learned query optimizers (LQOs) leverage historical execution data to enhance decision-making and generalize to unseen queries. In parallel, another line of research shifts part of the optimization responsibility into the execution phase, enabling dynamic plan adaptation based on actual runtime statistics[13, 14, 15, 16, 17, 18, 19, 20]. This paradigm, known as adaptive query processing (AQP), has been increasingly adopted by modern engines such as Spark SQL[21, 22], allowing certain decisions to be deferred until runtime.
While both approaches have significantly advanced the state of query optimization, they each face inherent limitations stemming from their core methodology. For LQOs, their optimization is confined to the pre-execution phase, which gives rise to two fundamental challenges:
C1: Lack of runtime information. Pre-execution query optimization requires the system to make planning decisions without access to essential runtime information, such as the actual cardinalities of intermediate results. Although researchers have proposed various encoding methods and predictive models to improve pre-execution estimation, these approaches still often fail to consistently outperform traditional rule-based optimizers such as PostgreSQL’s planner [23].
C2: Lack of fine-grained feedback. Most LQOs receive only a single final feedback metric (e.g., end-to-end execution time) for each query, even though they make multiple interdependent optimization decisions before execution. This coarse-grained signal fails to identify which specific decision led to performance changes. More critically, LQOs cannot observe intermediate execution states and revise their decisions based on actual runtime behavior.
On the other hand, prior work [24] has shown that even simple AQP techniques can, in some cases, match or even outperform more sophisticated LQOs. However, AQP methods also face two fundamental challenges:
C3: Lack of flexible adaptation. Most existing AQP methods rely on static, rule-based heuristics to trigger runtime adjustments, such as switching join strategies or enabling dynamic repartitioning. These rules are often tuned manually via parameters set by DBAs, which limits their flexibility and generalization to unseen workloads. Moreover, such methods rarely consider join reordering, despite it being one of the most impactful aspects of query optimization.
C4: Lack of practical deployability. Although AQP methods offer many valuable ideas, only a small number have been implemented in real-world database systems. A main barrier is the inherent complexity of combining planning and execution, which complicates testing, debugging, and performance tuning. Introducing learning-based mechanisms in this context further increases system complexity, making reliable deployment even more challenging.
This leads to a fundamental question: can we design a system that effectively combines the strengths of pre-execution learning with in-execution adaptation? Fig. 1 illustrates the design space of query optimizers along two key axes: when optimization occurs (pre-execution vs. in-execution) and how decisions are made (rule-based vs. learning-based). Most LQOs, such as AutoSteer [9] and Lero [10], perform the optimization in the pre-execution phase and make decisions without access to runtime feedback. In contrast, AQP methods, such as Spark SQL’s Adaptive Query Execution and runtime filters, perform limited rule-based adjustments during execution. As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned query optimizer that performs in-execution plan adaption on Spark SQL. It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision. It is designed as a plug-and-play module, requiring minimal changes to the underlying system, and can be easily integrated into Spark SQL.
Figure 1: AQORA is the first learned query optimizer achieving execution-time query plan optimization on Spark SQL.
AQORA addresses the aforementioned challenges through four key strategies:
S1: Runtime plan feature representation. AQORA removes the constraint of making all optimization decisions before execution begins. Instead, it dynamically collects runtime signals, such as actual cardinalities and execution statistics, and uses them as query plan features. This enables the model to make more informed and adaptive decisions based on real-time execution context. Compared to approaches that rely solely on pre-execution estimation [25, 26], this design also reduces the reliance on handcrafted features and complex estimators while improving robustness and accuracy.
[…]
S3: Learning-based adaptive strategies formulation. Instead of relying on fixed, rule-based heuristics, AQORA’s decision model automatically learns its optimization strategies from feedback collected during query execution. This data-driven approach enhances generalization across diverse query workloads without requiring manual tuning. To enhance learning stability and final performance, AQORA adopts a curriculum learning strategy: it starts with simpler decisions and gradually increases the complexity, enabling the model to learn more effective strategies in a step-by-step manner. In addition, AQORA supports a two-phase optimization mechanism, allowing strategies learned during execution to be reused in the planning phase of future queries.
S4: Lightweight and seamless system integration. Unlike prior approaches [10, 27] that require intrusive source code modifications, AQORA adopts a non-intrusive, plugin-style architecture. It can be enabled through simple configuration adjustments without altering the underlying query engine’s codebase. This plug-and-play design significantly reduces integration overhead and enables seamless deployment through minimal configuration.
Our contributions are summarized as follows:
  • We propose AQORA, an adaptive query optimizer for Spark SQL that integrates reinforcement learning into Spark’s adaptive query execution framework. AQORA comprises a decision model based on actor-critic learning and a planner extension that enables online plan modification during execution.
  • We adopt a lightweight decision model built upon the TreeCNN architecture [28], which learns optimization actions directly from stage-level feedback collected during query execution. By using true cardinalities as input features (and avoiding sample bitmaps or histogram encodings), the model achieves better query performance with a simplified architecture.
  • We design a plug-and-play planner extension built upon the extensibility interfaces of modern query engines, enabling non-intrusive integration with existing deployments. The extension intercepts runtime execution plans, interacts with the decision model, and applies optimization actions such as join reordering.
  • We conduct extensive experiments on three public benchmarks. Results show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
II Related Work
In this section, we review related work, including LQOs and previous AQP methods. Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
Plan-constructor
Plan-constructor methods aim to build query plans directly from scratch. These approaches enable the discovery of novel and potentially more efficient plans that lie beyond the capabilities of traditional optimizers. Neo [1] pioneers this category by leveraging past queries to progressively construct improved query plans, using the native optimizer as a weak oracle. Balsa [2] further improves the convergence efficiency of plan learning by employing enhanced initialization and training techniques. LOGER [3] enhances interpretability in learned planning by detecting detrimental operators and adjusting reward weighting accordingly. GLO [4] improves generalization to unseen workloads by clustering database statistics and learning generalized policies across these clusters. However, constructing plans from scratch suffers from two fundamental limitations: 1) the initial search often begins with poorly performing plans; and 2) the search space is extremely large, making efficient convergence challenging. In contrast, AQORA addresses these issues by starting from a reasonable initial plan generated by the native optimizer and learning to make incremental adjustments within a constrained action space.
Plan-steerer
Plan-steerer methods improve query plans by exploiting the internal mechanisms of the native optimizer. Instead of constructing plans from scratch, they guide the optimizer to explore multiple candidate plans and select the one with the lowest estimated cost. For example, Bao [5], DeepO [8], SCOPE [6], FASTgres [7], and AutoSteer [9] generate different query plans by toggling rule-based heuristics (i.e., hint sets) within traditional optimizers. HybridQO [11] explores alternative join orders by injecting leading hints. Lero[10] modifies cardinality estimates in the native optimizer to produce diverse plans. LEON [12] enumerates sub-plans directly using the native optimizer, while FOSS [30] refines plans after initial optimization by fixing join orders and methods. However, these methods rely heavily on the internal mechanisms of specific DBMS-native optimizers. As a result, the effectiveness of some of these methods is limited when applied to different DBMSes.
Adaptive query processing
AQP has been extensively studied to enable query plan adjustments during execution based on runtime feedback [13, 14, 15, 16, 18, 17, 19, 20, 27]. For example, Lookahead Information Passing [19] generalizes semi-join techniques to enhance the pipelining of equijoins. Adaptive Join Algorithm [31] dynamically switches between hash join and nested loop join during runtime. UDAO [27] adjusts parameters at multiple levels of granularity throughout query execution. RPT [32] extends predicate transfer with efficient semi-join reduction and join order validation to achieve provably robust join ordering for acyclic queries. In contrast to these methods, AQORA introduces the first learning-based AQP framework capable of continuously improving from experience and adaptively optimizing joins at runtime.
III Preliminary
In this section, we formulate the problem of query optimization and provide background on Spark SQL’s current approach.
III-A The Problem of Query Optimization
Query optimization involves many considerations, including predicate pushdown, expression simplification, and join optimization, etc. In this paper we focus on two core decisions: (1) determining the optimal join order and (2) selecting appropriate physical operators.
Consider a query $Q$ that joins $n$ tables $t_{1},t_{2},\cdots,t_{n}$}. If we restrict plan space to left-deep trees (Fig. 2(a)), there are up to $n!$ possible join orders. Allowing bushy trees (Fig. 2b) enlarges the search space to (2n2)!/(n1)!(2n-2)!/(n-1)! [33].
(a) left-deep join plan
(b) bushy join plan
Figure 2: In a left-deep tree plan, the left child of any join operation is a subtree, and the right child must be a leaf node. In contrast, bushy tree plans allow both children of any join operation to be subtrees
Join order and operator selection are critical factors in determining query execution efficiency. The core challenge of query optimization is to find a query plan $p$ that specifies both a join order and a set of physical operators, while minimizing the overall execution cost, which can be expressed as follows:
$\arg\min_{p}C(p)=C_{\text{plan}}(p)+C_{\text{execute}}(p)$
Here, $C_{\text{plan}}$ denotes the time spent generating candidate plans and identifying the final plan $p$, while $C_{\text{execute}}$ represents the actual execution time of plan $p$.
III-B Cost Based Optimization
Spark SQL’s cost-based optimizer (CBO) selects join orders using a dynamic programming algorithm that relies on cardinality estimates. This approach performs reasonably well for simple queries, where even with inaccurate cardinality estimates, suboptimal join orders still keep the planning cost ($C_{\text{plan}}$) within acceptable bounds. However, it encounters significant challenges when handling complex queries with many joins. The exponential time complexity of dynamic programming leads to rapid growth in optimization time, often resulting in both prolonged $C_{\text{plan}}$ and suboptimal plan selection. As shown in Fig. 3, $C_{\text{plan}}$ can become the dominant component of total cost as the number of joins increases. In fact, for JOB query 29a, the execution time shows little difference between enabling and disabling CBO, whereas the optimization time dominates when CBO is enabled. While we acknowledge that this issue may be specific to Spark SQL’s CBO implementation, our goal is not to replace or circumvent CBO but to provide a practical mechanism that dynamically mitigates such performance pitfalls.
Figure 3: Execution and optimization times of selected JOB queries with and without CBO.
CBO determines which join operator to use during physical optimization based on cardinality estimates. A key configuration is autoBroadcastJoinThreshold (BJT), which controls when broadcast joins are applied. If either side of the join is estimated to be smaller than BJT, Spark uses a broadcast hash join; otherwise, it falls back to shuffle joins such as sorted merge join.
However, this approach can lead to two problems. First, inaccurate cardinality estimates may result in selecting an inappropriate join type. Second, if the actual data size exceeds BJT but the estimate is incorrect, Spark may attempt to broadcast a large table, potentially causing worker crashes due to memory overload. To prevent such crashes, administrators often set BJT to a conservative value. While this safety measure enhances stability, it also limits the performance benefits offered by broadcast joins.
III-C Adaptive Query Execution
Spark SQL 3.0 (2020) introduced Adaptive Query Execution (AQE), which is an implementation of AQP and aims to tackle the challenges of cardinality estimation. As Xue et al. [22] claimed, maintaining accurate statistics in distributed systems like Spark is more difficult than in traditional databases because the continuous flow of data makes real-time statistics maintenance impractical. However, abandoning cardinality estimation is not feasible, as optimization without statistical guidance tends to perform even worse. AQE’s solution is to defer certain decisions (like determine join operators): it starts with rough estimates and then re-optimizes the query during execution using actual runtime metrics such as shuffle partition sizes, as illustrated in Fig. 4.
Figure 4: This query performs three joins: after the first join $t1\bowtie t2$ produces only 5 rows, AQE switches the second join from SMJ to BHJ; after the second join $(t1\bowtie t2)\bowtie t3$ produces 100,000 rows, AQE switches the third join from BHJ to SMJ.
AQE leverages runtime statistics to improve broadcast join decisions and adjust physical operators. It also performs specific optimizations such as merging shuffle partitions and fixing skewed joins. However, AQE cannot modify the initial join order determined during logical optimization and lacks the capability to learn from past executions.
IV System Overview
In this section, we introduce the main components of AQORA and its basic workflow.
As shown in Fig. 5, AQORA consists of two core components: a planner extension and a decision model. The workflow proceeds as follows:
Figure 5: An overview of AQORA, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
  1. 1.
    When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model. This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
  2. 2.
    The decision model analyzes the input plan and returns optimization actions (e.g., swapping the join order of leaf nodes).
  3. 3.
    The planner applies these actions to the ongoing query plan and resumes execution using the optimized version.
AQORA iteratively performs steps 1)-3) until the query completes and then performs step 4). We formalize this $k$-step process as a trajectory:
$\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$
where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding cardinalities $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$.
4) After the query completes, the decision model replays the collected trajectory to learn from past decisions and improve its optimization strategy for future queries.
V Decision Model
In this section, we describe the decision model in detail.
V-A The Actor-Critic Framework
We adopt the actor–critic framework to fully exploit the frequent feedback signals from AQE, which promotes faster convergence and allows us to observe the critic’s reward or penalty for each action taken by the actor.
V-A1 State, Action, and Return
To introduce the framework, we formally define the three main components of the Markov decision process for query optimization:
State
Each state $s_{t}=(p_{t},c_{t})$ consists of:
  • $p_{t}$: the partial executed plan at time step $t$;
  • $c_{t}$: the true cardinalities of completed plan segments, obtained from AQE statistics.
Action Space
AQORA supports four types of actions: turning on/off CBO, manipulating join orders, setting broadcast join hints on leaf nodes to enforce broadcast joins, and performing a no-op. The notations and definitions of these actions are summarized in Tab. I:
TABLE I: Action Space of AQORA’s Decision Model
Notation Definition
cbo(1/0)(1/0) turn on/off CBO
swap(i,j)(i,j) swap the $i$-th and $j$-th leaf node
[…]
broadcast(i)(i) broadcast the $i$-th leaf node
no-op do nothing
The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization. The first type (CBO action) is included because the quality of the initial query plan significantly impacts the subsequent optimization trajectory. Dynamically controlling this setting allows the agent to make better trade-offs between optimization quality and performance cost.
Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method. The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front.
For join method selection, the key issue is the overly conservative application of BJT. To address this, AQORA provides the broadcast action, which forces a broadcast join on a selected leaf node. While this strategy introduces some risk, it enables effective correction of conservative join method choices.
Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
Return
Based on empirical validation, we define the return as:
$R(\tau)=\sum_{i=1}^{k}\gamma^{i-1}r_{i}-\sqrt{T_{\text{execute}}(\tau)}$
where $T_{\text{execute}}$ represents the total execution time of $\tau$ (in seconds). The square root operation helps mitigate reward scale variations across different queries. If query execution fails, $R(\tau)$ is assigned a substantial negative penalty (e.g., $\sqrt{300}$). The intermediate reward $r_{i}$ is defined as the negative value of the additional number of shuffles introduced after taking action $a_{i}$, normalized by 10. This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process. The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $\sqrt{T_{\text{execute}}}$.
V-A2 Policy and State Transition
The actor network implements a stochastic policy $\pi_{\theta}(a|s)$, which maps states to action probabilities through softmax normalization:
$\pi_{\theta}(a|s)=\begin{cases}P_{1},\quad&\text{if}\ a=a_{1}\ (\text{e.g.,swap}\ (1,2))\\ P_{2},\quad&\text{if}\ a=a_{2}\ (\text{e.g.,broadcast}\ (1))\\ &\vdots\\ P_{m},\quad&\text{if}\ a=a_{m}\ (\text{e.g., no-op})\end{cases}$
where $\sum_{i=1}^{m}P_{i}=1$. The probabilities $P_{i}$ are computed by the actor network’s parameters $\theta$ via softmax normalization over its output logits. This probabilistic approach enables exploring alternative optimization paths while maintaining likely good choices.
After the actor takes an action $a$, the current state $s$ transitions to the next state $s^{\prime}$. The state transition probability $P(s^{\prime}|s,a)$ reflects the uncertainty in adaptive execution:
[…]
The uncertainty arises from AQE’s non-deterministic triggering: the execution engine may complete a varying number of stages between optimizations. As a result, applying the same action $a$ in state $s$ can lead to different successor states $s^{\prime}$.
V-A3 State Value and Action Value
The state value $v_{\pi}(s)$ quantifies the expected long-term utility of being in state $s$. The critic evaluates this value based on the current partial plan and runtime statistics. For example, if a join operation involves high cardinalities on both sides—often indicating expensive data shuffling—the critic may assign a lower $v_{\pi}(s)$ compared to states with smaller intermediate results. Note that this pattern isn’t universal—while some large joins may be unavoidable for optimal performance, it serves as an intuitive starting point. Through training, the critic learns to refine these initial heuristics into precise value estimates.
Also notice that the state value $v_{\pi}(s)$ depends not only on the current plan $s$ but also on the actor’s policy $\pi$. Even when reaching a promising state $s$, a suboptimal policy may select actions that degrade subsequent states. For example, a policy favoring broadcast joins for large tables could turn $s$ (with moderate cardinalities) into $s^{\prime}$ (with memory pressure), lowering $v_{\pi}(s)$ despite $s^{\prime}$’s inherent potential. The critic’s estimated value $v_{\phi}(s)$ approximates $v_{\pi}(s)$, learning to anticipate these policy-induced outcomes through exposure to the actor’s decision patterns.
The action value $q(s_{t},a_{t})$ measures the impact of action $a_{t}$ on optimization potential. Let $v_{\phi}(s)$ denote the critic’s value estimate with parameters $\phi$. We compute:
$q(s_{t},a_{t})=r_{t+1}+v_{\phi}(s_{t+1})-v_{\phi}(s_{t})$
where $s_{t+1}$ is the state after performing $a_{t}$. For instance, swapping larger tables early (action $a_{t}$) might decrease $v_{\phi}(s_{t+1})$ due to increased shuffle costs, yielding negative $q(s_{t},a_{t})$. Conversely, broadcasting a small table (action $a_{t}^{\prime}$) could improve $v_{\phi}(s_{t+1})$, producing positive action value. This differential signal guides the actor’s policy updates.
V-A4 Training Algorithm
For the actor, our goal is to maximize the expectation of $q(s_{t},a_{t})$ for each step. So we use the clipped PPO loss [34] with an entropy factor to encourage explorative actions:
$L^{\text{clip}}=-\frac{1}{k}\sum_{t=0}^{k-1}\left[\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}(a_{t}|s_{t})}},1\pm\epsilon\right)\right)q_{t}\right]$
$L^{\text{entropy}}=\frac{1}{k}\sum_{t=0}^{k-1}\pi_{\theta}(a_{t}|s_{t})\log\pi_{\theta}(a_{t}|s_{t})$
$L^{\text{actor}}=L^{\text{clip}}+\eta L^{\text{entropy}}$
For the critic, our goal is to minimize the bias of the evaluated state value $v_{\phi}(s)$ and the ground truth state value $v_{\pi}(s)$, so we use the MSE loss function to optimize the critic’s parameters $w$:
$L^{\text{critic}}=\frac{1}{k}\sum_{i=0}^{k-1}(v_{\phi}(s_{t})-v_{\pi}(s_{t}))^{2}$
We observe that the policy gradient for updating the actor depends on $q$ (calculated from the critic’s state values), while the critic’s update gradient relies on $v_{\pi}$ (determined by the actor’s current policy).
As shown in Alg. 1, this bootstrapping interaction allows both components to progressively improve through repeated optimization iterations, where $\gamma$ is set to 1 for simplicity.
Algorithm 1 The training algorithm in AQORA
1:Initial policy parameters $\theta$, initial value parameters $\phi$
2:Clip parameter $\epsilon$, training epochs $e$
3:Collect trajectory $(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$ using current policy $\pi_{\theta}$
4:Get actual state values $V_{\pi}:(v_{\pi}(s_{0}),v_{\pi}(s_{1}),\cdots v_{\pi}(s_{k}))$, $v_{\pi}(s_{i})=\sum_{j=i+1}^{k}r_{j}+\sqrt{T_{\text{execute}}}$
5:Get inferred state values $V_{\phi}:(v_{\phi}(s_{0}),v_{\phi}(s_{1}),\cdots v_{\phi}(s_{k}))$, $v_{\phi}(s_{i})$ is inferred from critic net’s initial parameters $\phi$
6:Compute action values $Q=(r_{1}+v_{\phi}(s_{1})-v_{\phi}(s_{0}),r_{2}+v_{\phi}(s_{2})-v_{\phi}(s_{1}),\cdots,r_{k}+v_{\phi}(s_{k})-v_{\phi}(s_{k-1}),0)$
7:Record old action probabilities $\pi_{\text{old}}=\pi_{\theta}(a_{t}|s_{t})$ inferred from actor net’s initial parameters $\theta$
8:for epoch =1=1 to $e$ do
9:  Compute new probabilities $\pi_{\theta}(a_{t}|s_{t})$
10:  Calculate probability ratio $r_{t}=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$
11:  Compute actor loss: $L^{\text{actor}}$
12:  Compute critic loss: $L^{\text{critic}}$
13:  Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}L^{\text{actor}}$
14:  Update $w\leftarrow w-\beta\nabla_{w}L^{\text{critic}}$
15:end for
From the above analysis, we see that the critic evaluates each individual action taken by the actor, providing feedback that is $k$ times denser than that of end-to-end reward methods. When the actor makes a suboptimal decision (e.g., selecting an inefficient join order), it receives immediate feedback via the critic’s updated state value, without needing to wait for the entire query to complete. Which means even if subsequent actions compensate for earlier mistakes (e.g., correcting a bad join order later in the plan), the initial suboptimal action still receives its own critical feedback. This ensures that every action contributes to policy updates, not just the final outcome.
V-B Model Design
This section presents the implementation details of the actor’s policy $\pi_{\theta}(s)$ and the critic’s value function $v_{\phi}(s)$.
When the decision model receives a partial plan $s$, it first compresses the plan tree and encodes each of its nodes into embedding vectors. The resulting vectorized tree is then processed separately by the actor and critic networks. The actor generates a probability distribution $\pi_{\theta}(s)$ over the set of valid actions, from which an action is selected using masked probabilities to filter out illegal operations. Meanwhile, the critic produces a scalar estimate $v_{\phi}(s)$, representing the state’s optimization potential.
V-B1 Tree Compressing
While nodes such as sorts and aggregates do affect query performance, we exclude them from the model’s input features to focus solely on join operations (Fig. 6(1)). This approach, though risky for other LQOs relying on full plan details, becomes viable because AQORA’s access to real-time cardinalities compensates for the lost information.
Figure 6: Tree compressing and feature encoding.
V-B2 Feature Encoding
The compressed plan tree is transformed into a vector tree, as shown in Fig. 6(2). Each node $u$ is encoded as a vector by concatenating three components:
[…]
with the following encoding rules:
  • Each node is assigned a type, such as a join or a broadcast query stage. AQORA encodes the node type as a one-hot vector type(u)(u). In a simplified example shown in Figure 6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
  • $\text{table}(u)$ is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Fig. 6. It encodes the join sequence leading to node $u$, and during AQE, even leaf nodes may touch multiple tables.
[…]
This encoding strategy ensures that AQORA remains operational under data updates and schema changes. Specifically, the operator type $\text{type}(u)$ is fixed, while the intermediate cardinality $\text{card}(u)$ is collected online during execution rather than relying on precomputed statistics. Even when new tables are introduced, the encoding remains valid, with the corresponding positions taking a default value of 0.
The encoded vector tree is processed by AQORA’s actor and critic networks.
We choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
V-B3 Masked Output
To prevent invalid actions such as swapping nonexistent nodes, we apply masking to the actor’s output. For a workload with at most $n$ tables, the dimension of the actor’s output is given by:
$d_{\pi_{\theta}(s)}=2+(n-1)+C_{n}^{2}+n+1$
for switching on/off CBO, n1n-1 possible leads (if we only consider lead a single table each time), $C_{n}^{2}$ possible swaps (each concrete action instance is treated as a distinct action, e.g., $\text{swap}(4,7)$), $n$ possible tables to broadcast, and 1 no-op. For example, when processing a partially executed plan with 3 leaf nodes, we create a binary mask as follows:
$\begin{split}M=(&\underbrace{0,0}_{\text{cbo(1), cbo(0), }},\underbrace{1,1,0,\cdots,1,0}_{\text{swap(1,2), swap(1,3), swap(2,3)}},\\ &\underbrace{1,1,0,\cdots}_{\text{lead(2), lead(3)}},\underbrace{1,1,1,0,\cdots}_{\text{broadcast(1), broadcast(2), broadcast(3)}},\underbrace{1}_{\text{no-op}})\end{split}$
However, to achieve smoother policy convergence, we implemented a curriculum learning framework that requires masking out additional actions at different training stages. Initially, the model is only allowed to choose whether to enable or disable CBO, focusing on a simple binary decision task. During the mid-training phase, we progressively lift the masking on runtime plan adjustment actions; as these decisions are based on actual runtime statistics, they are generally easier for the model to learn. In the final training stage, all restrictions are removed except for masking invalid actions, allowing the model to explore the full action space and refine its decision-making capabilities.
The final action probabilities are calculated by normalizing the masked output:
$\pi_{\text{final}}(s)=M\odot\pi_{\theta}(s)$
VI Planner Extension
Our planner extension operates through two core mechanisms: capturing and sending the current partial plan to the decision model and applying optimization actions returned by the actor to modify the query plan. As detailed at the end of Sec. VI-B1, these mechanisms support dynamic exploration of bushy joins through iterative plan modifications at runtime.
VI-A Send the Extracted Plan
The planner extension introduces a new optimization rule into existing rule set. This rule is triggered after all default (adaptive) optimization rules have been applied. Upon activation, it first checks whether the maximum number of allowed optimization steps has been reached. If not, it extracts the current partial execution plan along with available runtime cardinalities from each completed stage. The extracted plan and cardinalities are then sent to the decision model, after which the system awaits optimization instructions.
The setting of the maximum number of optimization steps plays a critical role in overall performance: higher values allow more optimization attempts and richer feedback signals, but they also require consistently accurate decisions across diverse scenarios. By default, we empirically set the maximum number of optimization steps (both before and during execution) to 3.
VI-B Inject Plan
The optimization instructions may involve five types of operations, as shown in Tab. I. While actions such as enabling CBO and no-op operations are straightforward to handle, different strategies are required for processing join reorder and broadcast instructions.
VI-B1 Plan Transformation
Swapping instructions are denoted as swap(i,j)(i,j), indicating the exchange of positions between the $i$-th and $j$-th leaf nodes. However, if the opposite sides of the join of these two leaf nodes lack compatible join keys, executing the swap would result in a Cartesian product. To mitigate this, we design Alg. 2 to perform controlled node swapping. Leading actions follow a similar strategy.
Algorithm 2 Swapping Algorithm
1:Plan $p$ with leaf nodes $l_{1},l_{2},\cdots,l_{n}]$ and join conditions $\{c_{1},c_{2},\dots,c_{m}\}$, swap instruction swap(i,j)(i,j) or lead instruction lead(i)(i)
2:New plan $p^{\prime}$ if the $i$-th and the $j$-th leaf nodes can be swapped or the $i$-th leaf node can be leaded, otherwise original plan $p$
[…]
5:Initialize $p^{\prime}\leftarrow l_{1}$
6:for $k=2$ to $n$ do
7:  Find condition $c\in C$ that cannot be evaluated separately on $p^{\prime}$ or $L^{\prime}_{k}$ but can be evaluated on $p^{\prime}\bowtie L^{\prime}_{k}$
8:  if $c$ exists then
9:   $p^{\prime}\leftarrow$ buildJoin($p^{\prime},L^{\prime}_{k},c$)
10:  else
11:   return $p$
12:  end if
13:end for
14:return $p^{\prime}$
We observe that applying swap (or lead) operations during a single execution can gradually transform the join plan into a bushy structure. As an illustrative example, consider the swap operation:
$((t_{1}\bowtie t_{2})\bowtie t_{3})\bowtie t_{4}\xrightarrow{\text{swap}((t_{1}\bowtie t_{2}),t_{4})}(t_{1}\bowtie t_{2})\bowtie(t_{3}\bowtie t_{4})$
This example highlights AQORA’s capability to explore bushy join plans through execution-time transformations. By applying the swap operation between the completed $t_{1}\bowtie t_{2}$ subtree and $t_{4}$, the system incrementally constructs a bushy plan during execution.
VI-B2 Join Operator Selection
Upon receiving a broadcast instruction, the planner extension performs a bottom-up traversal of the query plan. It first locates the corresponding join operator within the execution plan hierarchy, then annotates the identified join operator with a BROADCAST hint on the appropriate (left/right) side. This modification instructs Spark SQL to prioritize the broadcast join strategy for the specified join operation.
VII Experiments
In this section, we conduct extensive experiments to validate the effectiveness of AQORA.
VII-A Experiment Setup
VII-A1 Environment
All experiments were conducted on Spark 3.5.4 with 6 executors, each configured with 6 executor cores and 20 GB of executor memory. The decision model was trained and executed for inference on an NVIDIA A40 GPU.
VII-A2 Benchmarks
We selected three standard benchmarks, following prior work [4, 11].
JOB
[35] employs the real-world IMDb dataset with 21 tables related to movies and actors. It includes 113 complex SQL queries generated from 33 templates, each involving joins across 4 to 17 tables.
Extended JOB
[1] (referred to as ExtJOB) is an extension of the JOB benchmark, consisting of 24 queries derived from 12 distinct templates. Compared to the original JOB queries, ExtJOB features entirely different join graphs and predicates, making it a more challenging benchmark[2].
STACK
[5] comprises 10 tables from Stack Exchange datasets. Its 16 query templates generate analytical queries joining 4-12 relations. We excluded templates #4,7 (with arithmetic operations in join conditions) and #9,10 (containing unsupported predicate subqueries)—limitations shared by most existing LQOs. We plan to address these restrictions in the future.
VII-A3 Baselines
We compare AQORA against three state-of-the-art baselines:
Spark SQL
Combined with runtime filters and dynamic join selection, Spark SQL’s default configuration with AQE represents a strong baseline for adaptive query optimization, and it is also the most common industrial practice.
Lero
Lero [10] employs a learned comparator to evaluate join plans generated by traditional optimizers under different cardinality estimations, then chooses the best one among them.
AutoSteer
AutoSteer [9] systematically evaluates all available optimization rules in Spark SQL by disabling them to assess their impact on the current plan. It then constructs a collection of rules to disable for performance gains using greedy search.
We adopt the implementations of Lero and AutoSteer from their authors[36, 37]. We exclude other notable optimizers like Bao [5] and Balsa [2] primarily due to their incompatibility with Spark SQL, and we exclude UDAO[27] because it operates in a disjoint optimization space of parameter tuning.
VII-A4 Settings
Data Generation
For JOB and ExtJOB, we expanded the original IMDb dataset to ten times its original size. For STACK, we utilized the benchmark’s original archived dataset. All datasets were converted to Parquet format. Note that this experimental setup already imposes substantial memory pressure on our testing environment, as an bad query plan can easily lead to out-of-memory errors and execution failure.
Query Generation.
We generated 1,000 training queries using each benchmark’s templates. For each template, randomized predicate conditions were introduced while preserving the original join structure.
For JOB and ExtJOB, we exclusively used their original 113 and 24 queries as their test set, respectively. For STACK, the test set contains 10 additional queries per template generated in the same manner.
Evaluation Approach.
We conduct both static and dynamic evaluations. For the static evaluation, each optimizer is trained on the training set of a given benchmark until convergence, and then its performance is assessed by executing the corresponding test queries. For the dynamic evaluation, we train the optimizer on one dataset and then test it either on the test set of another dataset or on the same dataset after dynamically modifying the underlying data through updates.
Key Metrics
We evaluate performance using three core metrics: (1)End-to-end execution time $C$: total time required to run all test queries; (2)Optimization time $C_{\text{plan}}$: time spent generating optimization decisions; and (3)Raw execution time $C_{\text{execute}}$: time spent on purely executing the query plan. Each query’s execution time is capped at 300 seconds; queries that exceed this limit or fail due to out-of-memory are recorded as 300 seconds.
VII-B Query Performance
Fig. 7 shows the end-to-end execution time, optimization time, and raw execution time for all test queries across different methods.
Figure 7: Query performance on three benchmarks, in seconds. Black numbers outside bars represent end-to-end query execution time, while numbers in dark and light bars represent raw query time and optimization cost respectively.
VII-B1 End-to-End Execution Time
AQORA outperforms all three baselines across the three benchmarks. On the JOB benchmark, AQORA reduces end-to-end execution time by 47.3%, 90.1%, and 60.6% compared with Spark SQL’s default configuration, AutoSteer, and Lero, respectively. For ExtJOB, AQORA achieves reductions of 20.0%, 26.2%, and 38.5% against the same baselines. On STACK, AQORA reduces execution time by 73.3%, 17.2%, and 73.1% compared to Spark SQL’s default, AutoSteer, and Lero, respectively.
These results highlight the advantage of AQORA’s novel idea of optimizing queries during execution, surpassing traditional optimizers and other state-of-art methods.
VII-B2 Optimization Time
Spark SQL’s default optimization strategy incurs no additional overhead, as it directly executes the join order specified in the input SQL text. Consequently, we exclude its optimization time from our comparisons.
For learned methods such as Lero and AutoSteer, the optimization cost primarily stems from generating candidate plans. Lero achieves this by adjusting cardinality estimates at various levels of the query plan, producing over 10 candidate plans per query on average, with each EXPLAIN command taking about 10.1 seconds. AutoSteer, in contrast, generates more than 2 plans per query by enabling or disabling different optimization rules, and each EXPLAIN command takes around 3.3 seconds. This substantial overhead can, in some cases, lead to worse overall performance compared to Spark SQL’s default configuration, even when better query plans are discovered.
AQORA incurs significantly lower optimization costs, as it avoids the need to execute multiple EXPLAIN queries to enumerate candidate plans. Most of its overhead arises from model inference and the execution of Alg. 2, both of which are lightweight. However, AQORA introduces a new type of optimization overhead in distributed engines: modifying partially executed plans at runtime may trigger additional shuffle operations. Despite this, our experiments show the average optimization cost remains below 320 ms per query.
VII-B3 Raw Execution Time
The results show that the raw execution time of AQORA is comparable to that of Lero across all benchmarks. This indicates that modifying query plans during execution can be as effective as generating and evaluating multiple candidate plans from scratch.
The last row of Tab. II summarizes the number and proportion of inferior plans (i.e., out-of-memory or time-limit-exceeded) generated by each method across the three benchmarks. Spark SQL produces a substantial number of inferior plans, particularly on the STACK and JOB benchmarks, primarily due to its limited join optimization capabilities. AutoSteer also produces a number of inferior plans, as its learned optimization strategy tends to favor disabling high-overhead rules. While this can reduce planning time, it often backfires on complex queries. In contrast, AQORA does not generate any inferior plans during test query execution, although a few may arise during training. This highlights the stability and robustness of AQORA’s optimization strategy.
TABLE II: The number of queries showing performance improvement/regression compared with Spark SQL’s default configuration.
JOB ExtJOB STACK
$\delta\in$ Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA
$(0,0.2)$ - 3 1 41 - 4 0 9 - 19 8 22
$(0.2,+\infty)$ - 18 88 20 - 6 0 8 - 51 64 75
(0.2,0)(-0.2,0) - 15 11 41 - 9 8 5 - 19 29 7
$(-\infty,-0.2)$ - 77 13 11 - 5 16 2 - 21 9 6
Failure 10(8.8%) 0 20(17.6%) 0 0 0 0 0 33(30.0%) 0 34(30.1%) 0
VII-B4 Tail Latency
Tail performance is an important indicator of an optimizer’s stability. To provide a comprehensive view, we report the 30th, 60th, 90th, and 99th percentile execution times. Fig. 8 illustrates the end-to-end query latencies of Spark SQL, Lero, AutoSteer, and AQORA at these percentiles after training to convergence. We observe that AQORA consistently outperforms the other methods on all three benchmarks across most percentile levels.
Figure 8: Comparison of execution times at the 30th, 60th, 90th, and 99th percentiles of test queries across three benchmarks for different optimization methods.
On the JOB benchmark, AQORA achieves a 99th percentile latency of 66.86 s, yielding a 99.18% reduction compared to Lero. Notably, the other two baselines fail to produce executable plans at this percentile, underscoring AQORA’s robustness in handling complex queries and its significant advantage in tail latency reduction.
On the Stack benchmark, AQORA consistently outperforms all baselines at the 90th and 99th percentiles. Its raw execution latency at these percentiles is comparable to that of Lero, demonstrating that AQORA can achieve similarly high-quality plans without enumerating multiple candidates.
At lower percentiles, AQORA sometimes underperforms AutoSteer because the latter disables several costly optimization rules. For simpler queries that dominate these percentiles, skipping such rules yields better performance.
VII-B5 Dynamic Evaluation
The first row of Fig. 9 illustrates the performance of AQORA and Lero under data distribution changes. The full IMDb database we use contains data up to 2013. We construct two additional versions, IMDb-1950 and IMDb-1980, by filtering out all movies released after 1950 and 1980, along with their related entries in other tables. IMDb-1950 contains less than 10% of the full IMDb data, while IMDb-1980 retains about 30%[38]. We train the optimizers on these old versions of IMDb and then test them on the complete IMDb database to evaluate their robustness to data evolution. The second row of Fig. 9 shows the results where the optimizers are trained on the training set of JOB and tested on ExtJOB, and vice versa.
Figure 9: The first row shows the performance of AQORA and Lero when trained on IMDb-1950 and IMDb-1980 and tested on the full IMDb dataset. The second row presents the results where the optimizers are trained on the JOB and tested on ExtJOB, and vice versa.
We observe that AQORA consistently outperforms both Spark SQL and Lero under various experimental settings, except for a slight performance drop below Spark SQL when trained on ExtJOB queries using the IMDb-1950 dataset and tested on the full IMDb database. This result demonstrates AQORA’s robustness to dynamic environmental changes. Such robustness mainly stems from two factors: (1) the use of accurate runtime cardinalities enables AQORA to adapt its optimization behavior to the current execution context; and (2) the simplified feature encoding and lightweight network architecture effectively prevent overfitting.
However, both learning-based approaches experience noticeable performance degradation when the execution environment changes. AQORA shows relatively larger drops in raw execution performance. Specifically, Lero’s maximum raw execution degradation across the four settings is 10.80%, 7.86%, 7.53%, and 8.43%, while AQORA’s corresponding drops are 18.04%, 34.93%, 6.99%, and 18.33%. In many cases, AQORA fails to select the optimal plan because the required action was rarely explored during training, whereas Lero can almost always ensure that even an unchosen plan at training at least belongs to its explored set. If AQORA were trained using an offline RL algorithm equipped with a comprehensive experience buffer, this limitation might be mitigated.
VII-C Case Study
In this section, we analyze the results on a per-query basis to find out what AQORA actually achieves in optimizing query plans. In particular, we compare the end-to-end query execution time of learning-based methods with those generated by Spark SQL’s native optimizer. For a query plan $p$ executed by a learning-based method, we calculate its improvement or regression as $\delta=(C_{\text{Spark}}(p)-C_{\text{LQO}}(p))/C_{\text{Spark}}(p)$. Tab. II summarizes the distribution of $\delta$ for each method.
VII-C1 Query Improvement Analysis
As shown in Fig. 10, AQORA achieves up to 89.81%, 69.33%, and 91.83% performance improvement on individual queries from the JOB, ExtJOB, and STACK benchmarks, respectively.
Figure 10: The top 10 queries improved by AQORA in each benchmark compared with Spark SQL’s default configuration.
For example, in JOB query Q16a, AQORA not only enabled the CBO but also refined the plan by prioritizing the join between table title and table movie_keyword, significantly reducing the execution time from over 300 s to just 30.56 s, demonstrating AQORA’s ability to refine CBO-generated plans by improving join orders while balancing optimization overhead and performance gains.
[…]
VII-C2 Query Regression Analysis
While AQORA delivers substantial improvements on many queries, we also observe performance regressions on a small subset. Minor regressions are due to the optimization overhead introduced by AQORA.
Significant regressions stem from limitations of AQORA’s current learning framework, which occasionally leads to suboptimal plan choices (e.g., JOB Q28b). To mitigate this, we introduced several restrictions during training, such as limiting the number of candidate lead tables and capping the maximum number of action steps. These constraints help guide the agent toward more stable decisions and reduce exploration noise. However, such regressions are not entirely avoidable and remain an area for future improvement.
VII-C3 Proportion of bushy plans
Here we analyze the proportion of test queries whose final executed plans adopt a bushy join structure. In our experiments, 22 (19.5%) queries in JOB, 5 (20.1%) in ExtJOB, and 30 (27.3%) in STACK ultimately execute with bushy plans, enabled by the application of CBO actions and leading actions as described in Sec. VI-B1. These results demonstrate that AQORA is capable of effectively exploring bushy join plans to improve query performance.
VII-D Ablation Study
In this ablation study, we will try to replace AQORA’s core components with other possible substitutes to test out the effectiveness of AQORA’s design on ExtJOB. AQORA uses PPO as the default reinforcement learning algorithm, a TreeCNN architecture for both actor and critic and action networks, and an action space that includes CBO actions, leading actions (with only one allowed leaf node), and no-op actions. Learning is conducted through the curriculum process described in Sec.V-B3.
VII-D1 Different Reinforcement Learning Algorithms
To evaluate the impact of different reinforcement learning algorithms on query optimization performance, we compare the PPO algorithm used in AQORA with DQN [39]. As shown in Fig. 11(a), AQORA with PPO surpasses Spark SQL in execution latency significantly earlier during training compared to its DQN-based variant. Also, after training on 2400 queries, the PPO-based model achieves consistently lower execution latency than the DQN-based model. Although DQN benefits from experience replay, it suffers from limitations in environments with large action spaces and non-stationary dynamics.
Figure 11: Query performance on ExtJOB with different substituted modules, under varying numbers of training queries.
VII-D2 Different Neural Network Structures
We tested multiple neural network architectures for the actor and critic networks, with the basic model statistics and optimization overhead summarized in Tab. III. Among them, FCNN is the only model that fails to outperform Spark SQL even after training on 2400 queries. Specifically, LSTM exhibits the lowest optimization overhead (251 ms), making it more efficient during both training and inference. However, TreeCNN demonstrates better convergence behavior and achieves higher overall performance after training, albeit with a slightly increased overhead.
TABLE III: Parameters and optimization overhead of different model structure on JOB and ExtJOB dataset
Model Parameters Optimization overhead
TreeCNN 147506 317 ms per query
LSTM 4800 251 ms per query
QueryFormer 879602 671 ms per query
FCNN 410546 432 ms per query
VII-D3 Different Learning Strategy
To evaluate the effectiveness of our learning strategy design, we conducted an ablation study using two alternative strategies (Fig. 11(c)). In the first strategy, we removed the step limit during optimization while retaining the curriculum learning schedule. In the second strategy, we limited the maximum number of steps but removed the curriculum constraint, allowing AQORA to explore the full action space from the beginning.
Experimental results show that the first strategy performs comparably to the default configuration during the early training phase. However, as the curriculum progresses to its second stage, the model tends to take significantly more steps, making it difficult to repeatedly make correct decisions. The second strategy, despite enforcing a strict step limit, ultimately fails to outperform Spark SQL. This is because, in most cases, only a few targeted corrections are sufficient to reach an efficient plan. Without early-stage guidance, the model explores a large action space too freely, which hinders convergence.
VII-D4 Different Action Space
Although AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence. Specifically, we experimented with enabling or disabling the broadcast hint, the leading hint, and the CBO toggle. As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability.
When leading actions are disabled, AQORA loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries. Finally, disabling the CBO action prevents AQORA from jumping to an entirely different planning path starting from a CBO-generated plan. This restriction increases the difficulty for AQORA to escape local optima and discover globally better execution plans.
VIII Conclusion
We propose AQORA, the first reinforcement learning-based optimizer on Spark SQL that supports execution-time adaptation through four key innovations: (1) realistic feature encoding, (2) stage-level feedback with real-time intervention, (3) automatic strategy adaptation, and (4) seamless integration via Spark’s extension interface. By dynamically constructing bushy joins through node swapping and broadcast tuning, AQORA reduces Spark SQL’s execution latency by up to 70%, while achieving up to 99.8% faster optimization compared to enumeration-based learning query optimizers.

### v2

@@
-AQORA: A Learned Adaptive Query Optimizer for Spark SQL
+AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL
 Abstract
@@
 On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience.
-In this paper, we present AQORA, an adaptive query optimizer with a reinforcement learning architecture that combines the strengths of both LQO and AQP.
+In this paper, we present AQORA, a learned adaptive query optimizer that continuously refines query plans through stage-level feedback, unifying pre-execution and in-execution optimization within a single learning process.
 AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration.
@@
 Introducing learning-based mechanisms in this context further increases system complexity, making reliable deployment even more challenging.
-This leads to a fundamental question: can we design a system that effectively combines the strengths of pre-execution learning with in-execution adaptation?
+This leads to a fundamental question: can we design a query optimizer that continuously learns and adapts across both the pre- and in-execution phases?
 Fig.
@@
 In contrast, AQP methods, such as Spark SQL’s adaptive query execution and runtime filters, perform limited rule-based adjustments during execution.
-As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned query optimizer that performs in-execution plan adaption on Spark SQL.
+As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned adaptive query optimizer that performs in-execution plan refinement on Spark SQL.
 It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision.
@@
 II Related Work
-In this section, we review related work, including LQOs and previous AQP methods.
+In this section, we review related work on LQO and AQP.
 Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
@@
 In this section, we introduce the main components of AQORA and its basic workflow.
-As shown in Fig. 5, AQORA consists of two core components: a planner extension and a decision model. The workflow proceeds as follows:
+As shown in Fig.
+5, AQORA consists of two core components: a planner extension and a decision model.
+The figure illustrates an example where AQORA performs an optimization step during the execution of the STACK query q1#1.
 Figure 5: An overview of AQORA, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
@@
     When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model.
+In Fig.
+5, the system has just completed scanning table t and observed that the output cardinality is only 1.
+The planner extension then sends the partially executed plan and the row count (1) to the decision model.
 This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
   2. 2.
-    The decision model analyzes the input plan and returns optimization actions (e.g., swapping the join order of leaf nodes).
+    The decision model analyzes the input plan and returns optimization actions.
+As shown in Fig.
+5, the decision model observes that table t produces only one output row but is joined after two larger tables.
+The actor therefore infers that placing t earlier in the join order could be more efficient and outputs the action lead(t).
   3. 3.
-    The planner applies these actions to the ongoing query plan and resumes execution using the optimized version.
-AQORA iteratively performs steps 1)-3) until the query completes and then performs step 4). We formalize this $k$-step process as a trajectory:
+[…]
+4) After the query completes after $k$ iterations of 1)-3), the decision model replays the collected $k$ step optimization process to learn from past decisions and improve its optimization strategy for future queries.
+We formalize this optimization process as a trajectory:
 $\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$
 where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding cardinalities $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$.
-4) After the query completes, the decision model replays the collected trajectory to learn from past decisions and improve its optimization strategy for future queries.
 V Decision Model
@@
 The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization.
-The first type (CBO action) is included because the quality of the initial query plan significantly impacts the subsequent optimization trajectory.
-Dynamically controlling this setting allows the agent to make better trade-offs between optimization quality and performance cost.
+In some cases, enabling CBO produces a better starting plan that is closer to the optimal; in others, disabling it yields better results.
 Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method.
 The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front.
-For join method selection, the key issue is the overly conservative application of BJT.
-To address this, AQORA provides the broadcast action, which forces a broadcast join on a selected leaf node.
-While this strategy introduces some risk, it enables effective correction of conservative join method choices.
 Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
@@
 This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process.
-The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $\sqrt{T_{\text{execute}}}$.
+The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $R(\tau)$.
 V-A2 Policy and State Transition
@@
 AQORA encodes the node type as a one-hot vector type(u)(u).
-In a simplified example shown in Figure 6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
+In a simplified example shown in Fig.
+6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
   • $\text{table}(u)$ is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Fig.
@@
 Even when new tables are introduced, the encoding remains valid, with the corresponding positions taking a default value of 0.
-The encoded vector tree is processed by AQORA’s actor and critic networks.
-We choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
+The encoded vector tree is then processed by AQORA’s actor and critic networks, and we choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
 V-B3 Masked Output
@@
 VII-D4 Different Action Space
-Although AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence.
-Specifically, we experimented with enabling or disabling the broadcast hint, the leading hint, and the CBO toggle.
+AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence.
 As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability.
@@
 By dynamically constructing bushy joins through node swapping and broadcast tuning, AQORA reduces Spark SQL’s execution latency by up to 70%, while achieving up to 99.8% faster optimization compared to enumeration-based learning query optimizers.
+IX AI-Generated Content Acknowledgement
+Portions of this paper’s text were refined using ChatGPT 5 to improve the clarity, grammar, and conciseness of the language.
+The AI system was used only for linguistic polishing and did not contribute to the conceptual development, technical design, experimental results presented in this work.

### v3

@@
-AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL
-Abstract
-Recent studies have identified two main approaches to improve query optimization: learned query optimization (LQO), which generates or selects better query plans before execution based on models trained in advance, and adaptive query processing (AQP), which adapts the query plan during execution based on statistical feedback collected at runtime.
-Although both approaches have shown promise, they also face critical limitations.
-LQO must commit to a fixed plan without access to actual cardinalities and typically rely on a single end-to-end feedback signal, making learning inefficient.
-On the other hand, AQP depends heavily on rule-based heuristics and lacks the ability to learn from experience.
-In this paper, we present AQORA, a learned adaptive query optimizer that continuously refines query plans through stage-level feedback, unifying pre-execution and in-execution optimization within a single learning process.
-AQORA addresses the above challenges through four core strategies: (1) realistic feature encoding, (2) query stage-level feedback and intervention, (3) automatic strategy adaptation, and (4) low-cost integration.
-Experiments show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
-I Introduction
-Query optimization plays a central role in database systems, determining the physical plan for a given query.
-Typically, this process takes place prior to query execution.
-The query optimizer selects a fixed execution plan based on estimated data statistics, analytical cost models, and a predefined set of transformation rules.
-Once execution begins, the query processing engine adheres strictly to the chosen plan, without the opportunity to revise it based on actual runtime behavior.
-In response to the limitations of static, pre-execution query planning, subsequent research has evolved along two primary directions.
-One line of work focuses on developing learning-based algorithms that assist or replace the traditional query optimizer in making key planning decisions (such as join ordering, operator selection, or cardinality estimation) prior to execution beginning[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
-These learned query optimizers (LQOs) leverage historical execution data to enhance decision-making and generalize to unseen queries.
-In parallel, another line of research shifts part of the optimization responsibility into the execution phase, enabling dynamic plan adaptation based on actual runtime statistics[13, 14, 15, 16, 17, 18, 19, 20].
-This paradigm, known as adaptive query processing (AQP), has been increasingly adopted by modern engines such as Spark SQL[21, 22], allowing certain decisions to be deferred until runtime.
-While both approaches have significantly advanced the state of query optimization, they each face inherent limitations stemming from their core methodology.
-For LQOs, their optimization is confined to the pre-execution phase, which gives rise to two fundamental challenges:
-C1: Lack of runtime information.
-Pre-execution query optimization requires the system to make planning decisions without access to essential runtime information, such as the actual cardinalities of intermediate results.
-Although researchers have proposed various encoding methods and predictive models to improve pre-execution estimation, these approaches still often fail to consistently outperform traditional rule-based optimizers such as PostgreSQL’s planner [23].
-C2: Lack of fine-grained feedback.
-Most LQOs receive only a single final feedback metric (e.g., end-to-end execution time) for each query, even though they make multiple interdependent optimization decisions before execution.
-This coarse-grained signal fails to identify which specific decision led to performance changes.
-More critically, LQOs cannot observe intermediate execution states and revise their decisions based on actual runtime behavior.
-On the other hand, prior work [24] has shown that even simple AQP techniques can, in some cases, match or even outperform more sophisticated LQOs.
-However, AQP methods also face two fundamental challenges:
-C3: Lack of flexible adaptation.
-Most existing AQP methods rely on static, rule-based heuristics to trigger runtime adjustments, such as switching join strategies or enabling dynamic repartitioning.
-These rules are often tuned manually via parameters set by DBAs, which limits their flexibility and generalization to unseen workloads.
-Moreover, such methods rarely consider join reordering, despite it being one of the most impactful aspects of query optimization.
-C4: Lack of practical deployability.
-Although AQP methods offer many valuable ideas, only a small number have been implemented in real-world database systems.
-A main barrier is the inherent complexity of combining planning and execution, which complicates testing, debugging, and performance tuning.
-Introducing learning-based mechanisms in this context further increases system complexity, making reliable deployment even more challenging.
-This leads to a fundamental question: can we design a query optimizer that continuously learns and adapts across both the pre- and in-execution phases?
-Fig.
-1 illustrates the design space of query optimizers along two key axes: when optimization occurs (pre-execution vs.
-in-execution) and how decisions are made (rule-based vs.
-learning-based).
-Most LQOs, such as AutoSteer [9] and Lero [10], perform the optimization in the pre-execution phase and make decisions without access to runtime feedback.
-In contrast, AQP methods, such as Spark SQL’s adaptive query execution and runtime filters, perform limited rule-based adjustments during execution.
-As shown in the figure, our proposed system, AQORA (Adaptive Query Optimizer with Reinforcement learning Architecture), is the first to reside in the top-right quadrant: a learned adaptive query optimizer that performs in-execution plan refinement on Spark SQL.
-It integrates with Spark’s runtime engine via a lightweight planner extension and uses an actor-critic reinforcement learning model to make informed decisions during execution, leveraging query stage-level feedback for fine-grained supervision.
-It is designed as a plug-and-play module, requiring minimal changes to the underlying system, and can be easily integrated into Spark SQL.
-Figure 1: AQORA is the first learned query optimizer achieving execution-time query plan optimization on Spark SQL.
-AQORA addresses the aforementioned challenges through four key strategies:
-S1: Runtime plan feature representation.
-AQORA removes the constraint of making all optimization decisions before execution begins.
-Instead, it dynamically collects runtime signals, such as actual cardinalities and execution statistics, and uses them as query plan features.
-This enables the model to make more informed and adaptive decisions based on real-time execution context.
-Compared to approaches that rely solely on pre-execution estimation [25, 26], this design also reduces the reliance on handcrafted features and complex estimators while improving robustness and accuracy.
+LQRS: Learned Query Re-optimization Framework for Spark SQL
+Abstract.
+The query optimizer is a fundamental component of database management systems.
+Recent studies have shown that learned query optimizers outperform traditional cost-based query optimizers.
+However, they fail to exploit valuable runtime observations generated during query execution to dynamically re-optimize the plan, thereby limiting further improvements in query performance.
+To address this issue, we propose learned query re-optimization, which allows optimization decisions to be deferred to execution time and guided by actual runtime observations.
+We realize this idea through LQRS, a learned query re-optimization framework that builds upon Spark SQL, exploiting runtime observations for dynamic plan refinement.
+Specifically, LQRS employs a curriculum reinforcement learning strategy and jointly supports pre-execution and in-execution optimization, allowing knowledge learned during execution to directly benefit pre-execution planning.
+Furthermore, we design a plug-and-play planner extension built upon the extensibility interfaces of Spark SQL, enabling online plan modification.
+Experiments on Spark SQL demonstrate that LQRS reduces end-to-end execution time by up to 90% compared to other learned query optimizers and query re-optimization methods.
+PVLDB Reference Format:
+PVLDB, 14(1): XXX-XXX, 2020.
+doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
+doi:$X$.XX/XXX.XX
+PVLDB Artifact Availability:
+The source code, data, and/or other artifacts have been made available at https://github.com/heyewuyue1/LQRS.
+1. Introduction
+The query optimizer is a fundamental component of modern database management systems.
+Its primary role is to select an efficient execution plan for each query.
+Traditional cost-based query optimization (Selinger et al., 1979) is a common technique in query optimizers, where the idea is to select the plan with the minimum estimated cost.
+However, these cost models employ a variety of handcrafted formulas to approximate actual execution costs; such approximations are often inaccurate, resulting in suboptimal execution plans.
+To address this problem, people propose to employ machine learning in query optimization (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021; Chen et al., 2024b; Zhong et al., 2024; Sun et al., 2022; Yu et al., 2022; Chen et al., 2023b; Marcus et al., 2019; Yang et al., 2022; Chen et al., 2023a, 2024a).
+Such a learned query optimizer (LQO) leverages models trained on historical execution feedback to improve the enumeration and evaluation of candidate execution plans, thereby enabling more effective optimization decisions.
 […]
-S3: Learning-based adaptive strategies formulation.
-Instead of relying on fixed, rule-based heuristics, AQORA’s decision model automatically learns its optimization strategies from feedback collected during query execution.
-This data-driven approach enhances generalization across diverse query workloads without requiring manual tuning.
-To enhance learning stability and final performance, AQORA adopts a curriculum learning strategy: it starts with simpler decisions and gradually increases the complexity, enabling the model to learn more effective strategies in a step-by-step manner.
-In addition, AQORA supports a two-phase optimization mechanism, allowing strategies learned during execution to be reused in the planning phase of future queries.
-S4: Lightweight and seamless system integration.
-Unlike prior approaches [10, 27] that require intrusive source code modifications, AQORA adopts a non-intrusive, plugin-style architecture.
-It can be enabled through simple configuration adjustments without altering the underlying query engine’s codebase.
-This plug-and-play design significantly reduces integration overhead and enables seamless deployment through minimal configuration.
-Our contributions are summarized as follows:
-  • We propose AQORA, an adaptive query optimizer for Spark SQL that integrates reinforcement learning into Spark’s adaptive query execution framework.
-AQORA comprises a decision model based on actor-critic learning and a planner extension that enables online plan modification during execution.
-  • We adopt a lightweight decision model built upon the TreeCNN architecture [28], which learns optimization actions directly from stage-level feedback collected during query execution.
-By using true cardinalities as input features (and avoiding sample bitmaps or histogram encodings), the model achieves better query performance with a simplified architecture.
-  • We design a plug-and-play planner extension built upon the extensibility interfaces of modern query engines, enabling non-intrusive integration with existing deployments.
+Table 1.
+STACK Q1#1 involves optimizing a join plan over $t$, $s$, $q$, and tqtq.
+Lero fixes the join order before execution, whereas LQRS adjusts the join order based on runtime information, resulting in a shorter execution time.
+Plan Lero (Chen et al., 2024b) LQRS (ours)
+Pre-execution $(t\bowtie tq)\bowtie(s\bowtie q)$ $q\bowtie tq\bowtie t\bowtie s$
+In-execution Fixed during execution $t\bowtie tq\bowtie q\bowtie s*$
+Execution time 21.52s 15.24s
+    During execution, LQRS observes that the scan on table $t$ produces only one tuple and reorders the join accordingly.
+Such a limitation goes beyond LQO’s expressiveness or feature engineering.
+For example, characteristics such as the value distribution of aggregation keys or the exact selectivity of complex predicates only manifest once data is actually processed (Hellerstein et al., 1997).
+Even with sophisticated encodings and learned estimators, LQO models restricted to pre-execution optimization frequently struggle to produce reliable predictions for such properties, and empirical studies have shown that learned estimators often fail to consistently outperform traditional cost-based optimizers such as PostgreSQL’s planner (Heinrich et al., 2025).
+This fact suggests that the challenge arises not merely from modeling limitations, but from making decisions without access to critical runtime observations.
+This motivated us to think about an important question: how can we design a learned query re-optimization framework that can collect runtime observations and use them to dynamically re-optimize previously generated query plans during query execution?
+To address this issue, we propose learned query re-optimization, which allows optimization decisions to be deferred to execution time and guided by actual runtime observations.
+We realize this idea through LQRS, a learned query re-optimization framework that exploits runtime observations for dynamic plan refinement.
+Our realization of LQRS builds upon the Adaptive Query Execution (AQE) foundation of Spark SQL (Xue et al., 2024), which provides the necessary runtime hooks for collecting execution feedback and revising optimization decisions.
+As illustrated in Figure 1, existing query optimization techniques intervene at different phases of a query’s lifecycle: traditional rule-based operator optimization and cost-based join reordering are performed entirely before execution; most learned query optimizers such as Lero(Chen et al., 2024b) and AutoSteer (Anneser et al., 2023) are also restricted to the pre-execution phase; while adaptive query processing techniques like Spark AQE (Xue et al., 2024) and QuerySplit (Zhao et al., 2023) adjust execution without leveraging learning-based guidance.
+In contrast, LQRS represents the first learned query optimizer on Spark SQL that jointly supports pre-execution and in-execution plan refinement, effectively unifying learned query optimization with adaptive query processing.
+LQRS continuously collects runtime execution statistics and exposes them as plan features, including actual cardinality observed during query execution.
+These runtime features allow the optimizer to analyze partially executed plans using information unavailable before execution, enabling more effective optimization decisions during query processing.
+To enhance learning stability, LQRS adopts a curriculum learning strategy: it starts with simpler decisions and gradually increases the complexity, enabling the model to learn more effective strategies in a step-by-step manner.
+In addition, LQRS unifies pre-execution and in-execution plan optimization within a single learning process, allowing in-execution learned knowledge to be leveraged for pre-execution optimization.
+Figure 1. LQRS is the first learned query optimizer achieving both pre-execution and in-execution query plan optimization on Spark SQL.
+[…]
+Our contributions can be summarized as follows:
+  • We design and implement LQRS, a learned query re-optimizer which incorporates reinforcement learning into the Spark AQE framework to make decisions during query execution based on actual runtime observations.
+  • We adopt a decision model built upon the TreeCNN architecture  (Mou et al., 2016), which learns optimization actions directly from query stage-level feedback collected during query execution.
+By using true cardinalities and size in bytes as input features, the model achieves better query performance with a simplified architecture.
+  • We design a plug-and-play planner extension built upon the extensibility interfaces of Spark SQL, enabling online plan modification during execution.
 The extension intercepts runtime execution plans, interacts with the decision model, and applies optimization actions such as join reordering.
-  • We conduct extensive experiments on three public benchmarks.
-Results show that AQORA reduces end-to-end execution time by up to 90% compared to other learned methods and by up to 70% compared to Spark SQL’s default configuration with adaptive query execution.
-II Related Work
-In this section, we review related work on LQO and AQP.
-Since the pioneering work of LEO [29] in 2001, which introduced the concept of a learning-based query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
-Plan-constructor
+  • We conduct extensive experiments on four public benchmarks.
+Experiments on Spark SQL demonstrate that LQRS reduces end-to-end execution time by up to 90% compared to other learned query optimizers and query re-optimization methods.
+2. Related Work
+Figure 2.
+This query performs three joins: after the first join $t1\bowtie t2$ produces only 5 rows, AQE switches the second join from SMJ to BHJ; after the second join $(t1\bowtie t2)\bowtie t3$ produces 100,000 rows, AQE switches the third join from BHJ to SMJ.
+Learned Query Optimization
+Since LEO (Stillger et al., 2001) in 2001 introduced the concept of a learned query optimizer, recent studies on LQO have evolved into two main directions: plan-constructors and plan-steerers.
 Plan-constructor methods aim to build query plans directly from scratch.
-These approaches enable the discovery of novel and potentially more efficient plans that lie beyond the capabilities of traditional optimizers.
-Neo [1] pioneers this category by leveraging past queries to progressively construct improved query plans, using the native optimizer as a weak oracle.
-Balsa [2] further improves the convergence efficiency of plan learning by employing enhanced initialization and training techniques.
-LOGER [3] enhances interpretability in learned planning by detecting detrimental operators and adjusting reward weighting accordingly.
-GLO [4] improves generalization to unseen workloads by clustering database statistics and learning generalized policies across these clusters.
-However, constructing plans from scratch suffers from two fundamental limitations: 1) the initial search often begins with poorly performing plans; and 2) the search space is extremely large, making efficient convergence challenging.
-In contrast, AQORA addresses these issues by starting from a reasonable initial plan generated by the native optimizer and learning to make incremental adjustments within a constrained action space.
-Plan-steerer
+These approaches enable the discovery of novel and potentially more efficient plans.
+Neo  (Marcus et al., 2019) pioneers this category by leveraging past queries to progressively construct improved query plans, using the native optimizer as a weak oracle.
+Balsa  (Yang et al., 2022) further improves the convergence efficiency of plan learning by employing enhanced initialization and training techniques.
+LOGER  (Chen et al., 2023a) enhances interpretability in learned planning by detecting detrimental operators and adjusting reward weighting accordingly.
+GLO  (Chen et al., 2024a) improves generalization to unseen workloads by clustering database statistics and learning generalized policies across these clusters.
 Plan-steerer methods improve query plans by exploiting the internal mechanisms of the native optimizer.
 Instead of constructing plans from scratch, they guide the optimizer to explore multiple candidate plans and select the one with the lowest estimated cost.
-For example, Bao [5], DeepO [8], SCOPE [6], FASTgres [7], and AutoSteer [9] generate different query plans by toggling rule-based heuristics (i.e., hint sets) within traditional optimizers.
-HybridQO [11] explores alternative join orders by injecting leading hints.
-Lero[10] modifies cardinality estimates in the native optimizer to produce diverse plans.
-LEON [12] enumerates sub-plans directly using the native optimizer, while FOSS [30] refines plans after initial optimization by fixing join orders and methods.
-However, these methods rely heavily on the internal mechanisms of specific DBMS-native optimizers.
-As a result, the effectiveness of some of these methods is limited when applied to different DBMSes.
+For example, Bao  (Marcus et al., 2021), DeepO  (Sun et al., 2022), SCOPE  (Negi et al., 2021), FASTgres  (Woltmann et al., 2023), and AutoSteer  (Anneser et al., 2023) generate different query plans by toggling rule-based heuristics (i.e., hint sets) within traditional optimizers.
+HybridQO  (Yu et al., 2022) explores alternative join orders by injecting leading hints.
+Lero (Chen et al., 2024b) modifies cardinality estimates in the native optimizer to produce diverse plans.
+LQRS differs from existing LQOs in that it leverages runtime plan feature representations to make more informed optimization decisions and allows optimization decisions made pre-execution to be corrected during query execution.
 Adaptive Query Processing
-AQP has been extensively studied to enable query plan adjustments during execution based on runtime feedback [13, 14, 15, 16, 18, 17, 19, 20, 27].
-For example, Lookahead Information Passing [19] generalizes semi-join techniques to enhance the pipelining of equijoins.
-Adaptive Join Algorithm [31] dynamically switches between hash join and nested loop join during runtime.
-UDAO [27] adjusts parameters at multiple levels of granularity throughout query execution.
-RPT [32] extends predicate transfer with efficient semi-join reduction and join order validation to achieve provably robust join ordering for acyclic queries.
-In contrast to these methods, AQORA introduces the first learning-based AQP framework capable of continuously improving from experience and adaptively optimizing joins at runtime.
-III Preliminary
-In this section, we formulate the problem of query optimization and provide background on Spark SQL’s current approach.
-III-A The Problem of Query Optimization
-Query optimization involves many considerations, including predicate pushdown, expression simplification, and join optimization, etc.
-In this paper we focus on two core decisions: (1) determining the optimal join order and (2) selecting appropriate physical operators.
-Consider a query $Q$ that joins $n$ tables $t_{1},t_{2},\cdots,t_{n}$}.
-If we restrict plan space to left-deep trees (Fig.
-2(a)), there are up to $n!$ possible join orders.
-Allowing bushy trees (Fig.
-2b) enlarges the search space to (2n2)!/(n1)!(2n-2)!/(n-1)!
-[33].
-(a) left-deep join plan
-(b) bushy join plan
-Figure 2: In a left-deep tree plan, the left child of any join operation is a subtree, and the right child must be a leaf node.
-In contrast, bushy tree plans allow both children of any join operation to be subtrees
-Join order and operator selection are critical factors in determining query execution efficiency.
-The core challenge of query optimization is to find a query plan $p$ that specifies both a join order and a set of physical operators, while minimizing the overall execution cost, which can be expressed as follows:
-$\arg\min_{p}C(p)=C_{\text{plan}}(p)+C_{\text{execute}}(p)$
-Here, $C_{\text{plan}}$ denotes the time spent generating candidate plans and identifying the final plan $p$, while $C_{\text{execute}}$ represents the actual execution time of plan $p$.
-III-B Cost Based Optimization
-Spark SQL’s cost-based optimizer (CBO) selects join orders using a dynamic programming algorithm that relies on cardinality estimates.
-This approach performs reasonably well for simple queries, where even with inaccurate cardinality estimates, suboptimal join orders still keep the planning cost ($C_{\text{plan}}$) within acceptable bounds.
-However, it encounters significant challenges when handling complex queries with many joins.
-The exponential time complexity of dynamic programming leads to rapid growth in optimization time, often resulting in both prolonged $C_{\text{plan}}$ and suboptimal plan selection.
-As shown in Fig.
-3, $C_{\text{plan}}$ can become the dominant component of total cost as the number of joins increases.
-In fact, for JOB query 29a, the execution time shows little difference between enabling and disabling CBO, whereas the optimization time dominates when CBO is enabled.
-While we acknowledge that this issue may be specific to Spark SQL’s CBO implementation, our goal is not to replace or circumvent CBO but to provide a practical mechanism that dynamically mitigates such performance pitfalls.
-Figure 3: Execution and optimization times of selected JOB queries with and without CBO.
-CBO determines which join operator to use during physical optimization based on cardinality estimates.
-A key configuration is autoBroadcastJoinThreshold (BJT), which controls when broadcast joins are applied.
-If either side of the join is estimated to be smaller than BJT, Spark uses a broadcast hash join; otherwise, it falls back to shuffle joins such as sorted merge join.
-However, this approach can lead to two problems.
-First, inaccurate cardinality estimates may result in selecting an inappropriate join type.
-Second, if the actual data size exceeds BJT but the estimate is incorrect, Spark may attempt to broadcast a large table, potentially causing worker crashes due to memory overload.
-To prevent such crashes, administrators often set BJT to a conservative value.
-While this safety measure enhances stability, it also limits the performance benefits offered by broadcast joins.
-III-C Adaptive Query Execution
-Spark SQL 3.0 (2020) introduced adaptive query execution (AQE), which is an implementation of AQP and aims to tackle the challenges of cardinality estimation.
-As Xue et al.
-[22] claimed, maintaining accurate statistics in distributed systems like Spark is more difficult than in traditional databases because the continuous flow of data makes real-time statistics maintenance impractical.
-However, abandoning cardinality estimation is not feasible, as optimization without statistical guidance tends to perform even worse.
-AQE’s solution is to defer certain decisions (like determine join operators): it starts with rough estimates and then re-optimizes the query during execution using actual runtime metrics such as shuffle partition sizes, as illustrated in Fig.
-4.
-Figure 4: This query performs three joins: after the first join $t1\bowtie t2$ produces only 5 rows, AQE switches the second join from SMJ to BHJ; after the second join $(t1\bowtie t2)\bowtie t3$ produces 100,000 rows, AQE switches the third join from BHJ to SMJ.
-AQE leverages runtime statistics to improve broadcast join decisions and adjust physical operators.
-It also performs specific optimizations such as merging shuffle partitions and fixing skewed joins.
-However, AQE cannot modify the initial join order determined during logical optimization and lacks the capability to learn from past executions.
-IV System Overview
-In this section, we introduce the main components of AQORA and its basic workflow.
-As shown in Fig.
-5, AQORA consists of two core components: a planner extension and a decision model.
-The figure illustrates an example where AQORA performs an optimization step during the execution of the STACK query q1#1.
-Figure 5: An overview of AQORA, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
-  1. 1.
-    When Spark SQL’s AQE is triggered during execution, the planner extension extracts the current query plan along with actual cardinality statistics and forwards it to the decision model.
-In Fig.
-5, the system has just completed scanning table t and observed that the output cardinality is only 1.
+Adaptive query processing (AQP) has been extensively investigated to facilitate query plan adjustments during execution based on runtime feedback  (Amsaleg et al., 1996; Borovica-Gajic et al., 2015; Deshpande and Hellerstein, 2004; Deshpande et al., 2007; Kabra and DeWitt, 1998b; Eurviriyanukul et al., 2010; Zhu et al., 2017; Justen et al., 2024; Lyu et al., 2024).
+The primary AQP techniques are categorized as operator-based AQP and plan-based AQP.
+Operator-based AQP adapts the behavior of physical operators during execution.
+For instance, Lookahead Information Passing (Zhu et al., 2017) generalizes semi-join techniques to improve the pipelining of equijoins.
+The Adaptive Join Algorithm (Ward, 2019) dynamically selects between hash join and nested loop join during runtime.
+Plan-based AQP (or query re-optimization) re-invokes the optimizer at certain points during query execution to revise structural decisions such as join orders or the execution order of subqueries  (Kabra and DeWitt, 1998a; Markl et al., 2004; Neumann and Galindo-Legaria, 2013; Karanasos et al., 2014; Herlihy et al., 2024; Wang et al., 2023; Zhao et al., 2023).
+Following the pioneering work on dynamic re-optimization (Kabra and DeWitt, 1998a), Adaptive Metaprogramming (Herlihy et al., 2024) moves recursive query optimization and code generation from compile time to runtime, enabling customized join-order optimization based on runtime information.
+LPCE (Wang et al., 2023) adopts a two-stage learning strategy, consisting of an initial model and a refinement model, to progressively improve cardinality estimation accuracy for query operators.
+QuerySplit (Zhao et al., 2023) proposes a proactive query re-optimization framework that decomposes a query into multiple logical subqueries and determines their execution order according to estimated costs and output cardinalities.
+While most AQP techniques are not learning-based, LQRS extends the plan-based AQP paradigm and leverages reinforcement learning to guide optimization decisions based on runtime feedback.
+It also integrates pre-execution and in-execution optimization into a unified learning process and supports adaptation at both the operator and plan levels.
+3. Illustrative Overview
+3.1. Spark SQL’s Query Execution Model
+In Spark SQL, query execution is organized around query stages, which serve as the fundamental units of physical execution and runtime optimization.
+A query stage corresponds to a contiguous subgraph of the physical plan that can be executed without shuffle boundaries; stages are separated by shuffle operations that materialize intermediate results (for example, when a join is performed on a different join key from the preceding join, a shuffle is required to repartition the data, thereby introducing a new query stage).
+Once a query stage completes, Spark obtains accurate runtime statistics for that stage, including output cardinalities and data sizes, which are unavailable prior to execution.
+3.2. Spark SQL’s Adaptive Query Execution
+In order to tackle the challenges of cardinality estimation and physical plan cost estimation, Spark SQL 3.0 (2020) introduced AQE.
+AQE starts with an initial pre-execution query plan based on rough estimations and then re-optimizes it during execution using actual runtime observations.
+Figure 2 illustrates a mechanism adopted by AQE called dynamic join selection.
+In Figure 2, the query is executed in multiple query stages.
+The red segments in the figure correspond to standard query execution without adaptation, while the blue and green parts indicate the two plan modifications introduced by AQE interventions at runtime.
+After the first query stage, which executes the join $t1\bowtie t2$ and produces only 5 rows and 70 bytes of intermediate result, AQE decides to switch the join strategy for the second stage from SMJ to BHJ.
+After the second stage, which executes $(t1\bowtie t2)\bowtie t3$ and produces 100,000 rows and 700,000 bytes of intermediate result, AQE again adapts the plan, switching the join strategy for the third stage from BHJ back to SMJ.
+By reasoning at the granularity of query stages during execution, AQE can make more informed decisions based on actual intermediate cardinalities rather than relying solely on rough pre-execution estimates.
+3.3. Spark AQE’s Limitations
+Despite its ability to adapt join strategies at runtime, Spark AQE has inherent limitations.
+First, the effectiveness of AQE’s dynamic join selection is constrained by human-configured thresholds.
+In particular, AQE only converts an SMJ to a BHJ if the size of one side’s intermediate result is smaller than the broadcast join threshold (BJT).
+Since broadcasting a large table can lead to worker crashes, administrators typically configure BJT with an over-conservative static value, which limits the performance benefits.
+Second, AQE cannot change the join order once query execution begins.
+The optimizer relies on the pre-execution plan to determine the sequence of joins.
+As a result, queries with suboptimal join orders may still incur significant performance penalties, leaving untapped optimization potential that current AQE cannot address.
+Figure 3.
+An overview of LQRS, which consists of two main components: (1) a decision model that generates optimization actions, and (2) an AQE planner extension that applies these actions and provides feedback to the decision model.
+3.4. Our Solution
+LQRS is tightly integrated with Spark SQL’s query execution model.
+It treats the completion of a query stage as a natural decision point, where partially executed plans and query stage-level statistics are exposed to the decision model.
+By aligning optimization actions with query stage boundaries, LQRS ensures that runtime plan adaptations are both semantically valid and practically efficient within Spark SQL’s execution framework.
+Next, we introduce the main components of LQRS and its basic workflow.
+As shown in Figure 3, LQRS consists of two core components: a planner extension and a decision model.
+The figure illustrates an example where LQRS performs a single optimization action during the execution of the STACK query q1#1.
+  1. (1)
+    When Spark SQL’s AQE is triggered during execution, the planner extension extracts (①) the current query plan along with actual cardinality statistics and sends (②) it to the decision model.
+In Figure 3, the system has just completed scanning table t and observed the output cardinality.
 The planner extension then sends the partially executed plan and the row count (1) to the decision model.
 This process can also be triggered during the query optimization phase, in which case actual cardinalities are not available.
-  2. 2.
-    The decision model analyzes the input plan and returns optimization actions.
-As shown in Fig.
-5, the decision model observes that table t produces only one output row but is joined after two larger tables.
+  2. (2)
+    The decision model analyzes the input plan and returns (③) optimization actions.
+As shown in Figure 3, the decision model observes that table t produces only one output row but is joined after two larger tables.
 The actor therefore infers that placing t earlier in the join order could be more efficient and outputs the action lead(t).
-  3. 3.
-[…]
-4) After the query completes after $k$ iterations of 1)-3), the decision model replays the collected $k$ step optimization process to learn from past decisions and improve its optimization strategy for future queries.
-We formalize this optimization process as a trajectory:
-$\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$
-where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding cardinalities $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$.
-V Decision Model
-In this section, we describe the decision model in detail.
-V-A The Actor-Critic Framework
-We adopt the actor–critic framework to fully exploit the frequent feedback signals from AQE, which promotes faster convergence and allows us to observe the critic’s reward or penalty for each action taken by the actor.
-V-A1 State, Action, and Return
+  3. (3)
+    The planner extension applies (④) these actions to the ongoing query plan and injects (⑤) the optimized plan back to the AQE engine to resume execution.
+As illustrated in Figure 3, after receiving the action lead(t), the planner extension attempts to reposition table t lower in the plan, allowing it to participate in the join operation earlier.
+In coordination with the planner extension’s reposition behavior, the AQE engine can further adapt the join strategies, transforming the original SMJ–BHJ–BHJ sequence into a more efficient BHJ–SMJ–BHJ sequence.
+  4. (4)
+    After the query completes after $k$ iterations of ①-⑤, the decision model replays (⑥) the collected $k$ step optimization process to learn from past decisions and improve its optimization strategy for future queries.
+During this replay process, the critic evaluates (⑦) each action taken by the actor and provides feedback signals, which are used to guide policy improvement.
+We formalize the optimization process of LQRS as an episode:
+$\tau=(s_{0},a_{0},r_{1},s_{1},\cdots,a_{n-1},r_{k},s_{k})$
+where each state $s_{i}$ consists of a partially executed plan $p_{i}$ and its corresponding statistics $c_{i}$, while $a_{i}$ denotes the action taken at step $i$, and $r_{i+1}$ represents the shaping reward obtained after taking action $a_{i}$.
+Details will be given in the Section 4.1.1.
+3.5. Our Result
+As a result, compared to Spark AQE, LQRS reduces the runtime from 33.53s to 15.24s in the above example, achieving over 50% lower execution latency.
+In this example, Spark AQE is unable to reorder joins based on runtime observations and thus continues to execute the query using a suboptimal join order determined at pre-execution time, which both incurs higher execution cost and generates large intermediate results, preventing effective use of dynamic join selection.
+In contrast, LQRS captures runtime signals that reveal a better join order and proactively performs join reordering during execution, after which the AQE engine replans the remaining query with a more efficient join strategy.
+4. Decision Model
+4.1. The Actor-Critic Framework
+We adopt the actor–critic framework to fully exploit the frequent feedback signals from runtime, which promotes faster convergence and allows us to observe the critic’s reward or penalty for each action taken by the actor.
+4.1.1. State, Action, and Reward
 To introduce the framework, we formally define the three main components of the Markov decision process for query optimization:
-State
-Each state $s_{t}=(p_{t},c_{t})$ consists of:
-  • $p_{t}$: the partial executed plan at time step $t$;
-  • $c_{t}$: the true cardinalities of completed plan segments, obtained from AQE statistics.
-Action Space
-AQORA supports four types of actions: turning on/off CBO, manipulating join orders, setting broadcast join hints on leaf nodes to enforce broadcast joins, and performing a no-op.
-The notations and definitions of these actions are summarized in Tab.
-I:
-TABLE I: Action Space of AQORA’s Decision Model
+State. Each state $s_{t}=(p_{t},c_{t})$ consists of:
+  • $p_{t}$: the partially executed plan at time step $t$;
+  • $c_{t}$: the true cardinalities and output size in bytes of completed query stages, obtained from AQE statistics.
+Action Space.
+LQRS supports five actions: controlling the plan initialization strategy, swapping the $i$-th and $j$-th leaf node, joining the leaf nodes $t_{1},t_{2},\cdots,t_{l}$ first if possible, setting broadcast join hints on leaf nodes to enforce broadcast joins, and intentionally taking no action for now.
+The notations and definitions of these actions are summarized in Table 2.
+Table 2. Action space of LQRS’s decision model
 Notation Definition
-cbo(1/0)(1/0) turn on/off CBO
+init(1/0)(1/0) select plan initialization strategy
 swap(i,j)(i,j) swap the $i$-th and $j$-th leaf node
@@
 broadcast(i)(i) broadcast the $i$-th leaf node
-no-op do nothing
-The action space of AQORA is carefully designed to balance flexibility and tractability in adaptive query optimization.
-In some cases, enabling CBO produces a better starting plan that is closer to the optimal; in others, disabling it yields better results.
-Once an initial plan is chosen, AQORA can refine it through two kinds of structural modifications: adjusting the join order and modifying the join method.
-The lead action provides a more practical and efficient mechanism for guiding join order, as most beneficial changes can be achieved by repositioning a single leaf node to the front.
-Finally, the no-op action allows the model to intentionally defer decision and wait for more runtime information.
-Return
-Based on empirical validation, we define the return as:
-$R(\tau)=\sum_{i=1}^{k}\gamma^{i-1}r_{i}-\sqrt{T_{\text{execute}}(\tau)}$
-where $T_{\text{execute}}$ represents the total execution time of $\tau$ (in seconds).
-The square root operation helps mitigate reward scale variations across different queries.
-If query execution fails, $R(\tau)$ is assigned a substantial negative penalty (e.g., $\sqrt{300}$).
-The intermediate reward $r_{i}$ is defined as the negative value of the additional number of shuffles introduced after taking action $a_{i}$, normalized by 10.
-This definition naturally provides positive feedback when the model selects no-op (no additional shuffle is introduced), encouraging safe behavior during early training, while penalizing actions that introduce extra shuffles and discouraging costly changes late in the optimization process.
-The discount rate $\gamma(<1)$ ensures that the accumulated intermediate rewards $r_{i}$, even over many steps, do not dominate the overall cost $R(\tau)$.
-V-A2 Policy and State Transition
+no-op()() take no action for now
+  • The init action controls the initial plan construction strategy, determining how the starting execution plan is generated before any runtime adaptation takes place.
+In some cases, a traditional cost-based initialization produces a plan that is already close to optimal, while in others, a rule-based initialization offers a better starting point for subsequent optimization actions.
+Allowing the decision model to dynamically select between different plan initialization strategies helps LQRS to escape local optima and discover better execution plans.
+  • The swap action enables fine-grained join reordering by exchanging two leaf nodes in the current logical plan.
+This action provides local flexibility for adjusting join order and is particularly effective when small structural changes are sufficient to improve performance.
+  • The lead action promotes a specific table, or a subset of tables, to be joined earlier in the execution plan.
+Compared to the swap action, this action offers a more direct and efficient mechanism for correcting suboptimal join orders, as many beneficial transformations can be achieved by repositioning a single critical leaf node.
+As a result, the lead action serves as a practical control for guiding join order decisions.
+  • The broadcast action enforces a broadcast join on a selected leaf node.
+This action is designed to overcome the overly conservative behavior induced by static broadcast thresholds.
+Although forcing a broadcast join introduces some potential risks, it enables the correction of conservative join method choices.
+  • The no-op action allows the decision model to intentionally defer optimization decisions and wait for additional runtime information.
+This action is particularly useful in early execution stages where limited feedback is available and premature interventions may be harmful.
+By supporting deliberate inaction, LQRS can balance exploration and stability while reducing unnecessary plan perturbations.
+Together, these actions enable LQRS to perform both coarse-grained and fine-grained plan adaptations across different stages of query execution.
+Reward.
+A key challenge of runtime plan adaptation is that modifying an execution plan after some query stages have already materialized may require redistributing intermediate data.
+In Spark, this typically manifests as additional shuffle operations.
+For example, consider a query whose initial execution plan decides to evaluate $t_{1}\bowtie t_{2}$ first, followed by a join with $t_{3}$.
+During execution, Spark scans $t_{1}$ and $t_{2}$ and performs a shuffle to partition their records according to the join key of $t_{1}\bowtie t_{2}$, preparing for a shuffle-based join.
+If a subsequent runtime decision swaps the join order to instead evaluate $t_{1}\bowtie t_{3}$ first, the existing partitioning of $t_{1}$, which was produced to satisfy the join with $t_{2}$, is no longer compatible with the new join requirement.
+As a result, $t_{1}$ is reshuffled to match the join key of $t_{1}\bowtie t_{3}$, allowing execution to proceed correctly under the updated plan.
+Motivated by this observation, LQRS explicitly incorporates reshuffles as intermediate feedback in the reward definition, encouraging the agent to account for the operational cost of reorganizing intermediate data when making runtime decisions.
+Based on empirical validation, we define the total reward as:
+$R(\tau)=\sum_{i=1}^{k}\gamma^{i-1}r_{i}-\sqrt{T_{\text{execute}}(\tau)},$
+where $T_{\text{execute}}(\tau)$ denotes the total execution time of episode $\tau$ (in seconds).
+The square root operation is used to reduce reward scale variations across queries.
+If execution fails due to OOM errors or exceeds the time limit of 300s, $R(\tau)$ is assigned a substantial negative value (e.g., $\sqrt{300}$).
+The intermediate reward $r_{i}$ is defined as the negative number of additional shuffle operations introduced after taking action $a_{i}$, normalized by 10.
+This normalization aligns the scale of intermediate rewards with that of $\sqrt{T_{\text{execute}}(\tau)}$ under our experimental settings, ensuring that intermediate and terminal rewards are of comparable magnitude during training.
+This formulation naturally provides positive feedback when the agent selects a no-op action (i.e., no additional shuffle is introduced), encouraging conservative behavior in early training, while penalizing actions that trigger reshuffles and discouraging costly plan changes at later stages.
+The discount factor $\gamma<1$ further ensures that accumulated intermediate penalties do not dominate the overall reward.
+4.1.2. Policy and State Transition
 The actor network implements a stochastic policy $\pi_{\theta}(a|s)$, which maps states to action probabilities through softmax normalization:
-$\pi_{\theta}(a|s)=\begin{cases}P_{1},\quad&\text{if}\ a=a_{1}\ (\text{e.g.,swap}\ (1,2))\\ P_{2},\quad&\text{if}\ a=a_{2}\ (\text{e.g.,broadcast}\ (1))\\ &\vdots\\ P_{m},\quad&\text{if}\ a=a_{m}\ (\text{e.g., no-op})\end{cases}$
+$\pi_{\theta}(a|s)=\begin{cases}P_{1},\quad&\text{if}\ a=a_{1}\ (\text{e.g.,swap}(1,2))\\ P_{2},\quad&\text{if}\ a=a_{2}\ (\text{e.g.,broadcast}(1))\\ &\vdots\\ P_{m},\quad&\text{if}\ a=a_{m}\ (\text{e.g., no-op}())\end{cases}$
 where $\sum_{i=1}^{m}P_{i}=1$.
-The probabilities $P_{i}$ are computed by the actor network’s parameters $\theta$ via softmax normalization over its output logits.
+The probabilities $P_{i}$ are computed by the actor network with parameters $\theta$ via softmax normalization over its output logits.
 This probabilistic approach enables exploring alternative optimization paths while maintaining likely good choices.
@@
 […]
-The uncertainty arises from AQE’s non-deterministic triggering: the execution engine may complete a varying number of stages between optimizations.
+The uncertainty arises from Spark AQE’s non-deterministic triggering: the execution engine may complete a varying number of stages between optimizations.
 As a result, applying the same action $a$ in state $s$ can lead to different successor states $s^{\prime}$.
-V-A3 State Value and Action Value
+4.1.3. State Value and Action Value
 The state value $v_{\pi}(s)$ quantifies the expected long-term utility of being in state $s$.
@@
 This differential signal guides the actor’s policy updates.
-V-A4 Training Algorithm
-For the actor, our goal is to maximize the expectation of $q(s_{t},a_{t})$ for each step. So we use the clipped PPO loss [34] with an entropy factor to encourage explorative actions:
+4.1.4. Training Algorithm
+For the actor, our goal is to maximize the expectation of $q(s_{t},a_{t})$ for each step.
+So we use the clipped PPO loss (Schulman et al., 2017) with an entropy factor to encourage explorative actions:
 $L^{\text{clip}}=-\frac{1}{k}\sum_{t=0}^{k-1}\left[\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}(a_{t}|s_{t})}},1\pm\epsilon\right)\right)q_{t}\right]$
@@
 We observe that the policy gradient for updating the actor depends on $q$ (calculated from the critic’s state values), while the critic’s update gradient relies on $v_{\pi}$ (determined by the actor’s current policy).
-As shown in Alg. 1, this bootstrapping interaction allows both components to progressively improve through repeated optimization iterations, where $\gamma$ is set to 1 for simplicity.
-Algorithm 1 The training algorithm in AQORA
+As shown in Algorithm 1, this bootstrapping interaction allows both components to progressively improve through repeated optimization iterations, where $\gamma$ is set to 1 for simplicity.
+Algorithm 1 The training algorithm in LQRS
 1:Initial policy parameters $\theta$, initial value parameters $\phi$
 2:Clip parameter $\epsilon$, training epochs $e$
-3:Collect trajectory $(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$ using current policy $\pi_{\theta}$
+3:Collect episode $(s_{0},a_{0},r_{1},s_{1},\cdots,a_{k-1},r_{k},s_{k})$ using current policy $\pi_{\theta}$
 4:Get actual state values $V_{\pi}:(v_{\pi}(s_{0}),v_{\pi}(s_{1}),\cdots v_{\pi}(s_{k}))$, $v_{\pi}(s_{i})=\sum_{j=i+1}^{k}r_{j}+\sqrt{T_{\text{execute}}}$
@@
 13:  Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}L^{\text{actor}}$
-14:  Update $w\leftarrow w-\beta\nabla_{w}L^{\text{critic}}$
+14:  Update $\phi\leftarrow\phi-\beta\nabla_{\phi}L^{\text{critic}}$
 15:end for
@@
 This ensures that every action contributes to policy updates, not just the final outcome.
-V-B Model Design
+4.2. Model Design
 This section presents the implementation details of the actor’s policy $\pi_{\theta}(s)$ and the critic’s value function $v_{\phi}(s)$.
-When the decision model receives a partial plan $s$, it first compresses the plan tree and encodes each of its nodes into embedding vectors.
+When the decision model receives a partial plan $s$, it first encodes each of its nodes into embedding vectors.
 The resulting vectorized tree is then processed separately by the actor and critic networks.
-The actor generates a probability distribution $\pi_{\theta}(s)$ over the set of valid actions, from which an action is selected using masked probabilities to filter out illegal operations.
+The actor generates a probability distribution $\pi_{\theta}(s)$ over the set of valid actions.
 Meanwhile, the critic produces a scalar estimate $v_{\phi}(s)$, representing the state’s optimization potential.
-V-B1 Tree Compressing
-While nodes such as sorts and aggregates do affect query performance, we exclude them from the model’s input features to focus solely on join operations (Fig.
-6(1)).
-This approach, though risky for other LQOs relying on full plan details, becomes viable because AQORA’s access to real-time cardinalities compensates for the lost information.
-Figure 6: Tree compressing and feature encoding.
-V-B2 Feature Encoding
-The compressed plan tree is transformed into a vector tree, as shown in Fig. 6(2). Each node $u$ is encoded as a vector by concatenating three components:
+Figure 4. Feature encoding.
+4.2.1. Feature Encoding
+The plan tree is transformed into a vector tree, as shown in Figure 4. Each node $u$ is encoded as a vector by concatenating three components:
 […]
 with the following encoding rules:
-  • Each node is assigned a type, such as a join or a broadcast query stage.
-AQORA encodes the node type as a one-hot vector type(u)(u).
-In a simplified example shown in Fig.
-6, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0)
-  • $\text{table}(u)$ is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Fig.
-6.
+  • type(u)(u) is a one-hot vector encoding the type of node $u$, such as a join or a query stage.
+In the simplified example shown in Figure 4, if $u$ is a leaf node, then type(u)=(0,1)(u)=(0,1); if $u$ is a join node, then type(u)=(1,0)(u)=(1,0).
+  • table(u)(u) is a binary (0/1) vector indicating which tables participate in node $u$, as shown in Figure 4.
 It encodes the join sequence leading to node $u$, and during AQE, even leaf nodes may touch multiple tables.
-[…]
-This encoding strategy ensures that AQORA remains operational under data updates and schema changes.
+  • card(u)(u) is set to $(\log(1+c))$ if $u$ corresponds to a finished query stage with observed cardinality $c$; otherwise, card(u)(u) is set to (1)(-1).
+The same encoding strategy is applied to other statistics, such as size in bytes.
+This encoding strategy ensures that LQRS remains operational under data updates and schema changes.
 Specifically, the operator type $\text{type}(u)$ is fixed, while the intermediate cardinality $\text{card}(u)$ is collected online during execution rather than relying on precomputed statistics.
 Even when new tables are introduced, the encoding remains valid, with the corresponding positions taking a default value of 0.
-The encoded vector tree is then processed by AQORA’s actor and critic networks, and we choose TreeCNN for its low optimization overhead and superior performance after being trained on a sufficient number of queries (2400 on ExtJOB).
-V-B3 Masked Output
+The encoded vector tree is then processed by LQRS’s actor and critic networks, both of which are instantiated using a TreeCNN architecture.
+We refer the reader to  (Marcus et al., 2019) for a deeper investigation into tree convolution applied to query plans.
+4.2.2. Masked Output
 To prevent invalid actions such as swapping nonexistent nodes, we apply masking to the actor’s output. For a workload with at most $n$ tables, the dimension of the actor’s output is given by:
 $d_{\pi_{\theta}(s)}=2+(n-1)+C_{n}^{2}+n+1$
-for switching on/off CBO, n1n-1 possible leads (if we only consider lead a single table each time), $C_{n}^{2}$ possible swaps (each concrete action instance is treated as a distinct action, e.g., $\text{swap}(4,7)$), $n$ possible tables to broadcast, and 1 no-op.
+for 2 init actions, n1n-1 possible lead actions(if we only consider lead a single table each time), $C_{n}^{2}$ possible swap actions (each concrete action instance is treated as a distinct action, e.g., $\text{swap}(4,7)$), $n$ possible broadcast actions, and 1 no-op action.
 For example, when processing a partially executed plan with 3 leaf nodes, we create a binary mask as follows:
-$\begin{split}M=(&\underbrace{0,0}_{\text{cbo(1), cbo(0), }},\underbrace{1,1,0,\cdots,1,0}_{\text{swap(1,2), swap(1,3), swap(2,3)}},\\ &\underbrace{1,1,0,\cdots}_{\text{lead(2), lead(3)}},\underbrace{1,1,1,0,\cdots}_{\text{broadcast(1), broadcast(2), broadcast(3)}},\underbrace{1}_{\text{no-op}})\end{split}$
+$\begin{split}M=(&\underbrace{0,0}_{\text{init(1), init(0), }},\underbrace{1,1,0,\cdots,1,0}_{\text{swap(1,2), swap(1,3), swap(2,3)}},\\ &\underbrace{1,1,0,\cdots}_{\text{lead(2), lead(3)}},\underbrace{1,1,1,0,\cdots}_{\text{broadcast(1), broadcast(2), broadcast(3)}},\underbrace{1}_{\text{no-op}})\end{split}$
 However, to achieve smoother policy convergence, we implemented a curriculum learning framework that requires masking out additional actions at different training stages.
-Initially, the model is only allowed to choose whether to enable or disable CBO, focusing on a simple binary decision task.
+Initially, the model is only allowed to choose the plan initialization strategy, focusing on a simple binary decision task.
 During the mid-training phase, we progressively lift the masking on runtime plan adjustment actions; as these decisions are based on actual runtime statistics, they are generally easier for the model to learn.
@@
 $\pi_{\text{final}}(s)=M\odot\pi_{\theta}(s)$
-VI Planner Extension
-Our planner extension operates through two core mechanisms: capturing and sending the current partial plan to the decision model and applying optimization actions returned by the actor to modify the query plan.
-As detailed at the end of Sec.
-VI-B1, these mechanisms support dynamic exploration of bushy joins through iterative plan modifications at runtime.
-VI-A Send the Extracted Plan
-The planner extension introduces a new optimization rule into existing rule set.
-This rule is triggered after all default (adaptive) optimization rules have been applied.
-Upon activation, it first checks whether the maximum number of allowed optimization steps has been reached.
-If not, it extracts the current partial execution plan along with available runtime cardinalities from each completed stage.
-The extracted plan and cardinalities are then sent to the decision model, after which the system awaits optimization instructions.
+5. Planner Extension
+Our planner extension operates through two core mechanisms: extracting and sending the current partial plan to the decision model and applying optimization actions returned by the actor to modify the query plan.
+The planner extension is built on top of Spark AQE.
+In particular, the planner extension can rely on AQE’s native support for pausing query execution at well-defined synchronization points and triggering re-optimization based on runtime statistics.
+In Spark AQE, a series of built-in optimization rules (including dynamic join selection) is applied in a pipelined fashion each time AQE is triggered.
+Our planner extension is integrated as an extension rule of Spark AQE to implement five optimization actions, without requiring the planner extension to explicitly manage execution suspension or execution state preservation.
+5.1. Send the Extracted Plan
+Upon triggering the planner extension rule, it first checks whether the maximum number of allowed optimization steps has been reached.
+If not, it extracts the current partially executed plan along with available runtime statistics from each completed query stage.
+The extracted plan and statistics are then sent to the decision model, after which the planner extension awaits further actions.
 The setting of the maximum number of optimization steps plays a critical role in overall performance: higher values allow more optimization attempts and richer feedback signals, but they also require consistently accurate decisions across diverse scenarios.
 By default, we empirically set the maximum number of optimization steps (both before and during execution) to 3.
-VI-B Inject Plan
-The optimization instructions may involve five types of operations, as shown in Tab.
-I.
-While actions such as enabling CBO and no-op operations are straightforward to handle, different strategies are required for processing join reorder and broadcast instructions.
-VI-B1 Plan Transformation
-Swapping instructions are denoted as swap(i,j)(i,j), indicating the exchange of positions between the $i$-th and $j$-th leaf nodes.
+5.2. Inject Plan
+While init action and no-op action are straightforward to handle, different strategies are required for the swap action, lead action, and the broadcast action.
+5.2.1. Plan Transformation
+The swap action is denoted as swap(i,j)(i,j), indicating the exchange of positions between the $i$-th and $j$-th leaf nodes.
 However, if the opposite sides of the join of these two leaf nodes lack compatible join keys, executing the swap would result in a Cartesian product.
-To mitigate this, we design Alg.
-2 to perform controlled node swapping.
-Leading actions follow a similar strategy.
-Algorithm 2 Swapping Algorithm
-1:Plan $p$ with leaf nodes $l_{1},l_{2},\cdots,l_{n}]$ and join conditions $\{c_{1},c_{2},\dots,c_{m}\}$, swap instruction swap(i,j)(i,j) or lead instruction lead(i)(i)
+The same issue also occurs while applying the lead action.
+To mitigate this, we design Algorithm  2 to perform controlled plan transformation.
+Algorithm 2 The plan transformation algorithm in LQRS
+1:Plan $p$ with leaf nodes $l_{1},l_{2},\cdots,l_{n}]$ and join conditions $\{c_{1},c_{2},\dots,c_{m}\}$, swap action swap(i,j)(i,j) or lead action lead(i)(i)
 2:New plan $p^{\prime}$ if the $i$-th and the $j$-th leaf nodes can be swapped or the $i$-th leaf node can be leaded, otherwise original plan $p$
@@
 14:return $p^{\prime}$
-We observe that applying swap (or lead) operations during a single execution can gradually transform the join plan into a bushy structure. As an illustrative example, consider the swap operation:
-$((t_{1}\bowtie t_{2})\bowtie t_{3})\bowtie t_{4}\xrightarrow{\text{swap}((t_{1}\bowtie t_{2}),t_{4})}(t_{1}\bowtie t_{2})\bowtie(t_{3}\bowtie t_{4})$
-This example highlights AQORA’s capability to explore bushy join plans through execution-time transformations.
-By applying the swap operation between the completed $t_{1}\bowtie t_{2}$ subtree and $t_{4}$, the system incrementally constructs a bushy plan during execution.
-VI-B2 Join Operator Selection
-Upon receiving a broadcast instruction, the planner extension performs a bottom-up traversal of the query plan.
+We observe that applying swap or lead action during query execution can gradually transform the join plan into a bushy structure. As an illustrative example, consider the swap action in Figure  5:
+Figure 5. A swap between $(t_{1}\bowtie t_{2})$ and $t_{4}$ can turn the left-deep join plan into a bushy join plan.
+By applying the swap action between the completed $t_{1}\bowtie t_{2}$ subtree and $t_{4}$, the planner extension incrementally constructs a bushy plan during execution.
+5.2.2. Join Operator Selection
+Upon receiving a broadcast action, the planner extension performs a bottom-up traversal of the query plan.
 It first locates the corresponding join operator within the execution plan hierarchy, then annotates the identified join operator with a BROADCAST hint on the appropriate (left/right) side.
 This modification instructs Spark SQL to prioritize the broadcast join strategy for the specified join operation.
-VII Experiments
-In this section, we conduct extensive experiments to validate the effectiveness of AQORA.
-VII-A Experiment Setup
-VII-A1 Environment
-All experiments were conducted on Spark 3.5.4 with 6 executors, each configured with 6 executor cores and 20 GB of executor memory.
-The decision model was trained and executed for inference on an NVIDIA A40 GPU.
-VII-A2 Benchmarks
-We selected three standard benchmarks, following prior work [4, 11].
-JOB
-[35] employs the real-world IMDb dataset with 21 tables related to movies and actors. It includes 113 complex SQL queries generated from 33 templates, each involving joins across 4 to 17 tables.
-Extended JOB
-[1] (referred to as ExtJOB) is an extension of the JOB benchmark, consisting of 24 queries derived from 12 distinct templates.
-Compared to the original JOB queries, ExtJOB features entirely different join graphs and predicates, making it a more challenging benchmark[2].
-STACK
-[5] comprises 10 tables from Stack Exchange datasets.
+5.2.3. After Plan Injection
+After the optimization actions are applied, the modified plan is handed back to the AQE execution engine for continued execution.
+LQRS does not restart query execution or invalidate completed query stages.
+Instead, the injected modifications only affect the remaining unexecuted portion of the plan.
+In addition, the planner extension benefits from AQE’s built-in plan validation mechanisms.
+While LQRS does not rely on Spark’s internal plan evaluation or cost comparison and always adopts the plan produced by its optimization actions, AQE still enforces a set of critical safety checks before execution is resumed.
+In particular, the AQE engine guarantees the semantic equivalence of the modified plan with respect to the original one and ensures that the injected plan can be correctly compiled and executed by the underlying engine.
+As a result, the planner extension does not need to handle low-level concerns such as semantic equivalence checking, plan code generation, or execution resumption.
+5.3. Compatibility with Other Systems
+While the planner extension in LQRS is implemented on top of the Spark AQE framework, its design principles are applicable to other database systems.
+A key challenge in generalizing such extensions lies in identifying points at which execution can be paused or re-optimized with acceptable overhead.
+Spark SQL naturally provides such boundaries through query stages, which explicitly materialize intermediate results and support execution suspension and re-optimization.
+For database systems such as Flink SQL in batch mode, Presto, and Snowflake, which have similar execution models, natural synchronization points arise where intermediate results are produced.
+In these systems, planner extensions can likewise use these materialization boundaries to perform runtime optimization.
+For other systems that do not provide such exploitable execution boundaries, we need to identify points where execution can be safely paused or re-optimized.
+In PostgreSQL, We observe that pipeline-breaking operators, such as hash and sort operators, can be synchronization points because they inherently materialize intermediate results and interrupt tuple pipelines.
+DuckDB employs a vectorized execution engine, yet certain operators still require blocking behavior, creating implicit plan boundaries where runtime statistics can be observed.
+These observations indicate that even without explicit AQE support, our planner extension can use existing materialization or pipeline breakers for re-optimization.
+6. Experiments
+6.1. Experiment Setup
+6.1.1. Environment
+All experiments were conducted on Spark 3.5.4 with 6 executors, each configured with 6 cores and 20 GB memory. The decision model was trained and executed for inference on an NVIDIA A40 GPU.
+6.1.2. Benchmarks
+We selected four standard benchmarks, following prior work  (Chen et al., 2024a; Yu et al., 2022; Chen et al., 2024b):
+JOB (Leis et al., 2015): JOB employs the real-world IMDb dataset with 21 tables related to movies and actors.
+It includes 113 complex SQL queries generated from 33 templates, each of them involving joins across 4 to 17 tables.
+Extended JOB (referred to as ExtJOB(Marcus et al., 2019): ExtJOB is an extension of the JOB benchmark, consisting of 24 queries derived from 12 distinct templates.
+Compared to the original JOB queries, ExtJOB features entirely different join graphs and predicates, making it a more challenging benchmark (Yang et al., 2022).
+STACK (Marcus et al., 2021): STACK comprises 10 tables from Stack Exchange datasets.
 Its 16 query templates generate analytical queries joining 4-12 relations.
@@
 We plan to address these restrictions in the future.
-VII-A3 Baselines
-We compare AQORA against three state-of-the-art baselines:
-Spark SQL
-Combined with runtime filters and dynamic join selection, Spark SQL’s default configuration with AQE represents a strong baseline for adaptive query optimization, and it is also the most common industrial practice.
-Lero
-Lero [10] employs a learned comparator to evaluate join plans generated by traditional optimizers under different cardinality estimations, then chooses the best one among them.
-AutoSteer
-AutoSteer [9] systematically evaluates all available optimization rules in Spark SQL by disabling them to assess their impact on the current plan.
+TPC-H (1): TPC-H is designed to test the response time of database systems to complex queries.
+It includes eight tables with the scale factor set to 100 in our experiments.
+In our evaluation, we only consider the query templates #2, 3, 5, 7, 8, 9, 10 following  (Chen et al., 2024b).
+We exclude other templates, as they are too simple (only have one or two tables) or contain predicate subqueries.
+6.1.3. Baselines
+We compare LQRS against four baselines:
+Spark SQL: Spark SQL with its default AQE mechanism.
+Lero (Chen et al., 2024b): Lero employs a learned comparator to evaluate join plans generated by Spark SQL’s optimizer under different cardinality estimations, then chooses the best one among them.
+AutoSteer (Anneser et al., 2023): AutoSteer systematically evaluates all available optimization rules in Spark SQL by disabling them to assess their impact on the current plan.
 It then constructs a collection of rules to disable for performance gains using greedy search.
-We adopt the implementations of Lero and AutoSteer from their authors[36, 37].
-We exclude other notable optimizers like Bao [5] and Balsa [2] primarily due to their incompatibility with Spark SQL, and we exclude UDAO[27] because it operates in a disjoint optimization space of parameter tuning.
-VII-A4 Settings
-Data Generation
+SSA: QuerySplit (Zhao et al., 2023) is a query re-optimization algorithm that proactively decomposes a query into multiple subqueries at the logical plan level and iteratively selects subqueries to execute based on their estimated cost and output cardinality.
+In Spark SQL, query plans is naturally partitioned into query stages at shuffle boundaries.
+Therefore, we only re-implement the subquery selection component of QuerySplit, namely SSA, within Spark SQL and use it as a query re-optimization baseline for comparison.
+6.1.4. Settings
+For static evaluation, we refer to the experimental settings of Bao (Marcus et al., 2021) and Lero (Chen et al., 2024b), while for dynamic evaluation, we refer to the settings used in Robust-MSCN (Negi et al., 2023):
+Data Generation.
 For JOB and ExtJOB, we expanded the original IMDb dataset to ten times its original size.
 For STACK, we utilized the benchmark’s original archived dataset.
+For TPC-H, we generated the dataset using the official data generation tool with a scale factor of 100.
 All datasets were converted to Parquet format.
-Note that this experimental setup already imposes substantial memory pressure on our testing environment, as an bad query plan can easily lead to out-of-memory errors and execution failure.
+Note that this experimental setup already imposes substantial memory pressure on our testing environment, as a bad query plan can easily lead to out-of-memory errors and execution failure.
 Query Generation.
-We generated 1,000 training queries using each benchmark’s templates. For each template, randomized predicate conditions were introduced while preserving the original join structure.
+We generated 1,000 training queries using each benchmark’s templates.
+For each template, randomized predicate conditions were introduced while preserving the join structure.
 For JOB and ExtJOB, we exclusively used their original 113 and 24 queries as their test set, respectively.
-For STACK, the test set contains 10 additional queries per template generated in the same manner.
+For STACK and TPC-H, the test set contains 10 additional queries per template generated in the same manner.
 Evaluation Approach.
@@
 For the dynamic evaluation, we train the optimizer on one dataset and then test it either on the test set of another dataset or on the same dataset after dynamically modifying the underlying data through updates.
-Key Metrics
-We evaluate performance using three core metrics: (1)End-to-end execution time $C$: total time required to run all test queries; (2)Optimization time $C_{\text{plan}}$: time spent generating optimization decisions; and (3)Raw execution time $C_{\text{execute}}$: time spent on purely executing the query plan.
-Each query’s execution time is capped at 300 seconds; queries that exceed this limit or fail due to out-of-memory are recorded as 300 seconds.
-VII-B Query Performance
-Fig. 7 shows the end-to-end execution time, optimization time, and raw execution time for all test queries across different methods.
-Figure 7: Query performance on three benchmarks, in seconds.
-Black numbers outside bars represent end-to-end query execution time, while numbers in dark and light bars represent raw query time and optimization cost respectively.
-VII-B1 End-to-End Execution Time
-AQORA outperforms all three baselines across the three benchmarks.
-On the JOB benchmark, AQORA reduces end-to-end execution time by 47.3%, 90.1%, and 60.6% compared with Spark SQL’s default configuration, AutoSteer, and Lero, respectively.
-For ExtJOB, AQORA achieves reductions of 20.0%, 26.2%, and 38.5% against the same baselines.
-On STACK, AQORA reduces execution time by 73.3%, 17.2%, and 73.1% compared to Spark SQL’s default, AutoSteer, and Lero, respectively.
-These results highlight the advantage of AQORA’s novel idea of optimizing queries during execution, surpassing traditional optimizers and other state-of-art methods.
-VII-B2 Optimization Time
-Spark SQL’s default optimization strategy incurs no additional overhead, as it directly executes the join order specified in the input SQL text.
-Consequently, we exclude its optimization time from our comparisons.
-For learned methods such as Lero and AutoSteer, the optimization cost primarily stems from generating candidate plans.
-Lero achieves this by adjusting cardinality estimates at various levels of the query plan, producing over 10 candidate plans per query on average, with each EXPLAIN command taking about 10.1 seconds.
-AutoSteer, in contrast, generates more than 2 plans per query by enabling or disabling different optimization rules, and each EXPLAIN command takes around 3.3 seconds.
-This substantial overhead can, in some cases, lead to worse overall performance compared to Spark SQL’s default configuration, even when better query plans are discovered.
-AQORA incurs significantly lower optimization costs, as it avoids the need to execute multiple EXPLAIN queries to enumerate candidate plans.
-Most of its overhead arises from model inference and the execution of Alg.
-2, both of which are lightweight.
-However, AQORA introduces a new type of optimization overhead in distributed engines: modifying partially executed plans at runtime may trigger additional shuffle operations.
-Despite this, our experiments show the average optimization cost remains below 320 ms per query.
-VII-B3 Raw Execution Time
-The results show that the raw execution time of AQORA is comparable to that of Lero across all benchmarks.
-This indicates that modifying query plans during execution can be as effective as generating and evaluating multiple candidate plans from scratch.
-The last row of Tab.
-II summarizes the number and proportion of inferior plans (i.e., out-of-memory or time-limit-exceeded) generated by each method across the three benchmarks.
-Spark SQL produces a substantial number of inferior plans, particularly on the STACK and JOB benchmarks, primarily due to its limited join optimization capabilities.
-AutoSteer also produces a number of inferior plans, as its learned optimization strategy tends to favor disabling high-overhead rules.
-While this can reduce planning time, it often backfires on complex queries.
-In contrast, AQORA does not generate any inferior plans during test query execution, although a few may arise during training.
-This highlights the stability and robustness of AQORA’s optimization strategy.
-TABLE II: The number of queries showing performance improvement/regression compared with Spark SQL’s default configuration.
-JOB ExtJOB STACK
-$\delta\in$ Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA
-$(0,0.2)$ - 3 1 41 - 4 0 9 - 19 8 22
-$(0.2,+\infty)$ - 18 88 20 - 6 0 8 - 51 64 75
-(0.2,0)(-0.2,0) - 15 11 41 - 9 8 5 - 19 29 7
-$(-\infty,-0.2)$ - 77 13 11 - 5 16 2 - 21 9 6
-Failure 10(8.8%) 0 20(17.6%) 0 0 0 0 0 33(30.0%) 0 34(30.1%) 0
-VII-B4 Tail Latency
-Tail performance is an important indicator of an optimizer’s stability.
-To provide a comprehensive view, we report the 30th, 60th, 90th, and 99th percentile execution times.
-Fig.
-8 illustrates the end-to-end query latencies of Spark SQL, Lero, AutoSteer, and AQORA at these percentiles after training to convergence.
-We observe that AQORA consistently outperforms the other methods on all three benchmarks across most percentile levels.
-Figure 8: Comparison of execution times at the 30th, 60th, 90th, and 99th percentiles of test queries across three benchmarks for different optimization methods.
-On the JOB benchmark, AQORA achieves a 99th percentile latency of 66.86 s, yielding a 99.18% reduction compared to Lero.
-Notably, the other two baselines fail to produce executable plans at this percentile, underscoring AQORA’s robustness in handling complex queries and its significant advantage in tail latency reduction.
-On the Stack benchmark, AQORA consistently outperforms all baselines at the 90th and 99th percentiles.
-Its raw execution latency at these percentiles is comparable to that of Lero, demonstrating that AQORA can achieve similarly high-quality plans without enumerating multiple candidates.
-At lower percentiles, AQORA sometimes underperforms AutoSteer because the latter disables several costly optimization rules.
-For simpler queries that dominate these percentiles, skipping such rules yields better performance.
-VII-B5 Dynamic Evaluation
-The first row of Fig.
-9 illustrates the performance of AQORA and Lero under data distribution changes.
-The full IMDb database we use contains data up to 2013.
-We construct two additional versions, IMDb-1950 and IMDb-1980, by filtering out all movies released after 1950 and 1980, along with their related entries in other tables.
-IMDb-1950 contains less than 10% of the full IMDb data, while IMDb-1980 retains about 30%[38].
-We train the optimizers on these old versions of IMDb and then test them on the complete IMDb database to evaluate their robustness to data evolution.
-The second row of Fig.
-9 shows the results where the optimizers are trained on the training set of JOB and tested on ExtJOB, and vice versa.
-Figure 9: The first row shows the performance of AQORA and Lero when trained on IMDb-1950 and IMDb-1980 and tested on the full IMDb dataset.
-The second row presents the results where the optimizers are trained on the JOB and tested on ExtJOB, and vice versa.
-We observe that AQORA consistently outperforms both Spark SQL and Lero under various experimental settings, except for a slight performance drop below Spark SQL when trained on ExtJOB queries using the IMDb-1950 dataset and tested on the full IMDb database.
-This result demonstrates AQORA’s robustness to dynamic environmental changes.
-Such robustness mainly stems from two factors: (1) the use of accurate runtime cardinalities enables AQORA to adapt its optimization behavior to the current execution context; and (2) the simplified feature encoding and lightweight network architecture effectively prevent overfitting.
-However, both learning-based approaches experience noticeable performance degradation when the execution environment changes.
-AQORA shows relatively larger drops in raw execution performance.
-Specifically, Lero’s maximum raw execution degradation across the four settings is 10.80%, 7.86%, 7.53%, and 8.43%, while AQORA’s corresponding drops are 18.04%, 34.93%, 6.99%, and 18.33%.
-In many cases, AQORA fails to select the optimal plan because the required action was rarely explored during training, whereas Lero can almost always ensure that even an unchosen plan at training at least belongs to its explored set.
-If AQORA were trained using an offline RL algorithm equipped with a comprehensive experience buffer, this limitation might be mitigated.
-VII-C Case Study
-In this section, we analyze the results on a per-query basis to find out what AQORA actually achieves in optimizing query plans.
-In particular, we compare the end-to-end query execution time of learning-based methods with those generated by Spark SQL’s native optimizer.
-For a query plan $p$ executed by a learning-based method, we calculate its improvement or regression as $\delta=(C_{\text{Spark}}(p)-C_{\text{LQO}}(p))/C_{\text{Spark}}(p)$.
-Tab.
-II summarizes the distribution of $\delta$ for each method.
-VII-C1 Query Improvement Analysis
-As shown in Fig. 10, AQORA achieves up to 89.81%, 69.33%, and 91.83% performance improvement on individual queries from the JOB, ExtJOB, and STACK benchmarks, respectively.
-Figure 10: The top 10 queries improved by AQORA in each benchmark compared with Spark SQL’s default configuration.
-For example, in JOB query Q16a, AQORA not only enabled the CBO but also refined the plan by prioritizing the join between table title and table movie_keyword, significantly reducing the execution time from over 300 s to just 30.56 s, demonstrating AQORA’s ability to refine CBO-generated plans by improving join orders while balancing optimization overhead and performance gains.
+Key Metrics.
+We evaluate performance using End-to-end execution time: total time required to run all test queries.
+Each query’s end-to-end execution time is capped at 300s; queries that exceed this limit or fail due to OOM error are recorded as 300s.
+6.2. Query Performance
+Table 3 reports the end-to-end execution time across all benchmarks.
+Table 3. Overall performance on all benchmarks.
+Time (s) JOB ExtJOB STACK TPC-H
+Spark SQL 5821.4 694.0 13149.1 2954.0
+Lero 33211.9 752.9 4231.6 2625.0
+AutoSteer 7793.3 903.2 13026.1 3010.2
+SSA 3124.1 651.0 7934.2 4895.7
+LQRS 3069.5 555.5 3504.3 2471.3
+LQRS consistently achieves the best performance on all four workloads, outperforming all baseline methods.
+Compared with Spark SQL, LQRS reduces the total execution time by 16.3%–73.3% across four benchmarks.
+These results highlight the advantage of LQRS’s novel idea of allowing learned optimization decisions to be deferred to execution time and guided by actual runtime observations.
+Figure 6 illustrates the performance improvement and regression of each query on four tested benchmarks with Lero, AutoSteer, SSA, and LQRS w.r.t the corresponding end-to-end query execution latency in Spark SQL.
+We omit cases where Spark SQL chooses a similar plan (with a runtime difference << 5s) to LQRS and other baseline methods.
+The key observations are as follows:
+Figure 6. Latency improvement and regression of each query in Lero, AutoSteer, SSA, and LQRS w.r.t Spark SQL
+  • Compared with Lero, our performance improvements exhibit similar trends, but we observe fewer regressions, especially on JOB.
+This indicates that LQRS can leverage runtime observations to obtain execution plans of comparable quality to Lero, while avoiding the extensive plan enumeration and evaluation overhead that Lero incurs, highlighting the effectiveness of LQRS.
+  • Compared with AutoSteer, our improvements are substantially larger.
+AutoSteer’s learned optimization strategy tends to favor disabling high-overhead rules, which often backfires on complex queries.
+In contrast, LQRS does not generate any inferior plans during test query execution, highlighting the robustness of its optimization strategy.
+  • Compared with SSA, our performance gains are more substantial and stable, with fewer large-scale regressions observed on TPC-H.
+This is because our reinforcement learning–based decision process is more comprehensive: it evaluates feedback at each step of the actor and considers the long-term value of partially executed plans.
+As a result, LQRS does not greedily lead low-cost leaf nodes forward, but instead balances immediate costs with long-term benefits, highlighting the intelligence of its decision-making process.
+6.3. Case Study
+In this section, we analyze the results on a per-query basis to find out what LQRS actually achieves in optimizing query plans.
+6.3.1. Query Improvement Analysis
+Figure  7 shows the top 10 queries improved by LQRS in each benchmark compared with Spark SQL.
+LQRS achieves up to 89.81%, 69.33%, 91.83%, and 55.21% performance improvement on the JOB, ExtJOB, STACK, and TPC-H benchmarks, respectively.
+Figure 7. The top 10 queries improved by LQRS in each benchmark compared with Spark SQL.
+For example, in JOB query Q16a, LQRS not only chooses the cost-based initialization strategy but also refines the plan by prioritizing the join between table t and table mk, significantly reducing the execution time from over 300s to just 30.56s, demonstrating LQRS’s ability to refine CBO-generated plans by improving join orders while balancing optimization overhead and performance gains.
 […]
-VII-C2 Query Regression Analysis
-While AQORA delivers substantial improvements on many queries, we also observe performance regressions on a small subset. Minor regressions are due to the optimization overhead introduced by AQORA.
-Significant regressions stem from limitations of AQORA’s current learning framework, which occasionally leads to suboptimal plan choices (e.g., JOB Q28b).
-To mitigate this, we introduced several restrictions during training, such as limiting the number of candidate lead tables and capping the maximum number of action steps.
-These constraints help guide the agent toward more stable decisions and reduce exploration noise.
-However, such regressions are not entirely avoidable and remain an area for future improvement.
-VII-C3 Proportion of bushy plans
-Here we analyze the proportion of test queries whose final executed plans adopt a bushy join structure.
-In our experiments, 22 (19.5%) queries in JOB, 5 (20.1%) in ExtJOB, and 30 (27.3%) in STACK ultimately execute with bushy plans, enabled by the application of CBO actions and leading actions as described in Sec.
-VI-B1.
-These results demonstrate that AQORA is capable of effectively exploring bushy join plans to improve query performance.
-VII-D Ablation Study
-In this ablation study, we will try to replace AQORA’s core components with other possible substitutes to test out the effectiveness of AQORA’s design on ExtJOB.
-AQORA uses PPO as the default reinforcement learning algorithm, a TreeCNN architecture for both actor and critic and action networks, and an action space that includes CBO actions, leading actions (with only one allowed leaf node), and no-op actions.
-Learning is conducted through the curriculum process described in Sec.V-B3.
-VII-D1 Different Reinforcement Learning Algorithms
-To evaluate the impact of different reinforcement learning algorithms on query optimization performance, we compare the PPO algorithm used in AQORA with DQN [39].
-As shown in Fig.
-11(a), AQORA with PPO surpasses Spark SQL in execution latency significantly earlier during training compared to its DQN-based variant.
+6.3.2. Query Regression Analysis
+While LQRS delivers substantial improvements on many queries, we also observe performance regressions on a small subset.
+Minor regressions are mainly due to the optimization overhead introduced by LQRS.
+Significant regressions in query performance mostly stem from inherent limitations of LQRS’s current learning framework, which occasionally results in suboptimal plan selections.
+For instance, JOB Q8d experiences performance degradation because the table t was incorrectly selected as the lead table, causing a substantial increase in intermediate data movement.
+As a result, the shuffle size during query execution nearly tripled compared to the baseline plan, significantly impacting runtime.
+To mitigate such regressions, we introduced several training-time constraints, including limiting the number of candidate lead tables and capping the maximum number of action steps.
+These restrictions guide the agent toward more stable and predictable decisions while reducing exploration noise.
+Nevertheless, regressions of this kind cannot be entirely avoided under the current framework and represent an important avenue for future improvement.
+6.3.3. Proportion of bushy plans
+Here, we analyze the proportion of test queries whose final executed plans adopt a bushy join structure.
+In our experiments, 22 (19.5%) queries in JOB, 5 (20.1%) in ExtJOB, and 30 (27.3%) in STACK ultimately execute with bushy plans, enabled by the plan transformation actions as described in Section 5.2.1.
+These results show that LQRS is capable of effectively exploring bushy join plans to improve query performance.
+6.4. Dynamic Evaluation
+Figure 8 evaluates the robustness of LQRS and Lero under data distribution and workload changes. We chose Lero due to its relatively strong adaptability among all the baselines.
+For data distribution changes (Figure 8(a–b)), we construct two historical variants of IMDb—IMDb-1950 and IMDb-1980—by retaining only movies released no later than 1950 and 1980, respectively, together with their related records.
+These variants contain less than 10% and about 30% of the full IMDb dataset (Negi et al., 2023).
+The optimizers are trained on these historical datasets and evaluated on the full IMDb to assess robustness to distribution shifts.
+For workload changes (Figure 8(c–d)), the optimizers are trained on either JOB or ExtJOB and evaluated on both workloads, enabling cross-workload generalization analysis.
+Figure 8.
+Performance of LQRS and Lero under data distribution and workload changes.
+(a–b) Optimizers trained on older IMDb snapshots (1950, 1980) and tested on the full database.
+(c–d) Optimizers trained on either JOB or ExtJOB workloads and tested on both JOB and ExtJOB.
+We observe that LQRS consistently outperforms both Spark SQL and Lero under various experimental settings, except for a slight performance drop below Spark SQL when trained on ExtJOB queries using the IMDb-1950 dataset and tested on the full IMDb database.
+This result demonstrates LQRS’s robustness to environmental changes.
+Such robustness mainly stems from two factors: (1) the use of runtime observation enables LQRS to adapt its optimization behavior to the current execution context; and (2) the simplified feature encoding architecture effectively prevents overfitting.
+However, both learning-based approaches suffer performance degradation under execution environment changes, with LQRS exhibiting relatively larger drops in raw execution performance.
+Across the four settings, Lero’s maximum degradation ranges from 7.53% to 10.80%, whereas LQRS experiences more severe drops of up to 34.93%.
+In many cases, LQRS fails to select the optimal plan because the required action was rarely explored during training, whereas Lero can almost always ensure that even an unchosen plan at training at least belongs to its explored set.
+If LQRS were trained using an offline RL algorithm equipped with a comprehensive experience buffer, this limitation might be mitigated.
+6.5. Ablation Study
+In this ablation study, we will try to replace LQRS’s core components with other possible substitutes to test out the effectiveness of LQRS’s design on ExtJOB.
+We choose ExtJOB because it is relatively small, making experiments more convenient to run, yet it is still sufficient to reveal performance trends.
+LQRS uses PPO as the default reinforcement learning algorithm, a TreeCNN architecture for both actor and critic and action networks, and an action space that includes init action, lead action (with only one allowed leaf node), and no-op action.
+Learning is conducted through the curriculum process described in Section 4.2.2.
+6.5.1. Different Reinforcement Learning Algorithms
+To evaluate the impact of different reinforcement learning algorithms on query optimization performance, we compare the PPO algorithm used in LQRS with DQN  (Mnih et al., 2015).
+As shown in Figure 9(a), LQRS with PPO surpasses Spark SQL in execution latency significantly earlier during training compared to its DQN-based variant.
 Also, after training on 2400 queries, the PPO-based model achieves consistently lower execution latency than the DQN-based model.
 Although DQN benefits from experience replay, it suffers from limitations in environments with large action spaces and non-stationary dynamics.
-Figure 11: Query performance on ExtJOB with different substituted modules, under varying numbers of training queries.
-VII-D2 Different Neural Network Structures
-We tested multiple neural network architectures for the actor and critic networks, with the basic model statistics and optimization overhead summarized in Tab.
-III.
+Figure 9. Query performance on ExtJOB with different substituted modules, under varying numbers of training queries.
+6.5.2. Different Neural Network Structures
+We experimented with several model architectures proposed in prior work, including a fully connected neural network (FCNN), TreeCNN  (Marcus et al., 2021), LSTM  (Zhu et al., 2015), and QueryFormer  (Zhao et al., 2022).
+To adapt these models for LQRS, we replaced their original feature encoding with our own and substituted their scalar output layers with a LogSoftmax layer.
 Among them, FCNN is the only model that fails to outperform Spark SQL even after training on 2400 queries.
-Specifically, LSTM exhibits the lowest optimization overhead (251 ms), making it more efficient during both training and inference.
+Specifically, LSTM exhibits the lowest optimization overhead, making it more efficient during both training and inference.
 However, TreeCNN demonstrates better convergence behavior and achieves higher overall performance after training, albeit with a slightly increased overhead.
-TABLE III: Parameters and optimization overhead of different model structure on JOB and ExtJOB dataset
-Model Parameters Optimization overhead
-TreeCNN 147506 317 ms per query
-LSTM 4800 251 ms per query
-QueryFormer 879602 671 ms per query
-FCNN 410546 432 ms per query
-VII-D3 Different Learning Strategy
-To evaluate the effectiveness of our learning strategy design, we conducted an ablation study using two alternative strategies (Fig.
-11(c)).
+6.5.3. Different Learning Strategy
+To evaluate the effectiveness of our learning strategy design, we conducted an ablation study using two alternative strategies (Figure 9(c)).
 In the first strategy, we removed the step limit during optimization while retaining the curriculum learning schedule.
-In the second strategy, we limited the maximum number of steps but removed the curriculum constraint, allowing AQORA to explore the full action space from the beginning.
+In the second strategy, we limited the maximum number of steps but removed the curriculum constraint, allowing LQRS to explore the full action space from the beginning.
 Experimental results show that the first strategy performs comparably to the default configuration during the early training phase.
@@
 This is because, in most cases, only a few targeted corrections are sufficient to reach an efficient plan.
-Without early-stage guidance, the model explores a large action space too freely, which hinders convergence.
-VII-D4 Different Action Space
-AQORA’s planner extension supports a wide range of action types, we conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence.
-As shown in the results, enabling the broadcast hint often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability.
-When leading actions are disabled, AQORA loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries.
-Finally, disabling the CBO action prevents AQORA from jumping to an entirely different planning path starting from a CBO-generated plan.
-This restriction increases the difficulty for AQORA to escape local optima and discover globally better execution plans.
-VIII Conclusion
-We propose AQORA, the first reinforcement learning-based optimizer on Spark SQL that supports execution-time adaptation through four key innovations: (1) realistic feature encoding, (2) stage-level feedback with real-time intervention, (3) automatic strategy adaptation, and (4) seamless integration via Spark’s extension interface.
-By dynamically constructing bushy joins through node swapping and broadcast tuning, AQORA reduces Spark SQL’s execution latency by up to 70%, while achieving up to 99.8% faster optimization compared to enumeration-based learning query optimizers.
-IX AI-Generated Content Acknowledgement
-Portions of this paper’s text were refined using ChatGPT 5 to improve the clarity, grammar, and conciseness of the language.
-The AI system was used only for linguistic polishing and did not contribute to the conceptual development, technical design, experimental results presented in this work.
+Without early-stage guidance, the model explores an overly large action space too freely, which hinders convergence.
+6.5.4. Different Action Space
+LQRS’s planner extension supports a wide range of action types.
+We conducted ablation studies to identify which subsets of the action space contribute most to learning efficiency and convergence.
+As shown in the results, enabling the broadcast action often leads the model to generate plans that broadcast excessively large tables during training, resulting in significant performance fluctuations and instability.
+When the lead action is disabled, LQRS loses its ability to explicitly manipulate join ordering, thereby reducing its optimization power and leading to suboptimal plans for many complex queries.
+Finally, disabling the init action prevents LQRS from jumping to an entirely different planning path starting from a CBO-generated plan.
+This restriction increases the difficulty for LQRS to escape local optima and discover better execution plans.
+6.6. Overhead Study
+Table 4. Optimization cost on all benchmarks.
+Time (s) JOB ExtJOB STACK TPC-H
+Lero 30098.1 201.3 1110.6 379.5
+AutoSteer 829.2 176.7 634.0 442.1
+SSA 81.5 6.2 41.0 27.7
+LQRS 33.3 7.1 37.3 19.6
+As shown in Table 4, LQRS achieves the lowest optimization overhead among most evaluated methods, by making re-optimization decisions during query execution based on accurate runtime information.
+Its design explicitly supports a no-op action, allowing the optimizer to skip unnecessary plan modifications when the expected benefit is limited, which effectively suppresses redundant re-optimizations and keeps the optimization process lightweight.
+In comparison, SSA incurs higher overhead because it continuously modifies the plan whenever the current leading leaf node is not estimated as the lowest-cost node, resulting in unnecessary plan modifications during execution.
+Learned query optimizers such as Lero and AutoSteer suffer from even greater costs due to explicit enumeration and evaluation of a large plan search space before execution.
+Lero perturbs cardinality estimates at multiple levels to generate numerous alternative plans, while AutoSteer explores different combinations of optimization rules, both requiring repeated cost estimation over multiple candidates.
+7. Conclusion
+We propose LQRS, a learned query re-optimizer integrated into Spark SQL.
+LQRS serves as a prototype system that demonstrates how execution-time feedback and intervention can be incorporated into a modern query engine via a planner extension.
+By continuously observing partial execution results and applying query stage-level feedback with runtime-aware rewards, LQRS incrementally refines the remaining execution plan during query execution.
+Extensive experiments on Spark SQL demonstrate that LQRS reduces end-to-end execution time by up to 90% compared to other learned query optimizers and query re-optimization methods.